### v1

# Gap-preserving reductions and RE-completeness of independent set games

Abstract
In complexity theory, gap-preserving reductions play a crucial role in studying hardness of approximation and in analyzing the relative complexity of multiprover interactive proof systems. In the quantum setting, multiprover interactive proof systems with entangled provers correspond to gapped promise problems for nonlocal games, and the recent result MIP=RE [JNV+20a] shows that these are in general undecidable. However, the relative complexity of problems within MIP is still not well-understood, as establishing gap-preserving reductions in the quantum setting presents new challenges.
In this paper, we introduce a framework to study such reductions and use it to establish MIP-completeness of the gapped promise problem for the natural class of independent set games. In such a game, the goal is to determine whether a given graph contains an independent set of a specified size. We construct families of independent set games with constant question size for which the gapped promise problem is undecidable. In contrast, the same problem is decidable in polynomial time in the classical setting.
To carry out our reduction, we establish a new stability theorem, which could be of independent interest, allowing us to perturb families of almost PVMs to genuine PVMs.
1 Introduction
Interactive proof systems are a central concept in computational complexity theory. Shortly after the introduction of the multi-prover variant, Babai, Fortnow, and Lund characterized its power by showing that MIP = NEXP [BFL91], where MIP is the class of languages that can be decided by a multi-prover interactive proof system and NEXP is the set of languages with exponential-time checkable proofs. In fact, already MIP(2,1) = NEXP i.e. only two provers and a single round of interaction suffice to reach NEXP [FL92]. Such single-round proof systems can be viewed as families of nonlocal games $C=\{\mathcal{G}_{x}\}_{x}$, where the all-powerful provers attempt to convince the efficient verifier that the word $x\in\{0,1\}^{*}$ belongs to a language $L$, see Section 2.2 for a definition of nonlocal games. The power of MIP(2,1) is captured by the gapped promise problem for the family of games $C$: for fixed constants $0<s<c\leq 1$, given a game $\mathcal{G}_{x}\in C$, decide whether the value of $\mathcal{G}_{x}$ is at least $c$, or less than $s$, given the promise that one of the two must hold. We denote this as the $(c,s)$-gap problem for $C$ and call $c$ and $s$ the completeness and soundness parameter respectively. Importantly, the families of games considered in the above results are succinctly presented, i.e., while the question and answer sets can grow exponentially in $\left|x\right|$, there are efficient Turing machines $S$ and $V$ to respectively sample questions and verify the answers from the provers, see Section 2.2 for a formal definition. In fact, the PCP theorem – the corner stone for the classical theory of hardness of approximation – can be viewed as a scaled-down version of MIP = NEXP, where the families of games have question and answer sets which grow at most polynomially in $\left|x\right|$ [ALM+98, AS98].
Since the provers are assumed to be all-powerful, it is natural to allow them to take advantage of quantum physics and most notably shared entanglement. The complexity class MIP, where the provers can share entanglement, is captured by the gapped promise problem defined above where we replace the value of a game by its quantum value [CHTW04]. We will refer to this as the $(c,s)$-gap problem. In comparison to MIP, its quantum counterpart MIP proved much harder to understand. Over the years, MIP was shown to contain MIP and increasingly larger complexity classes [IV12, Vid16, Vid20, Ji16, Ji17, NV18a, NV18b, FJVY19, NW19]. These developments were further re-enforced by a result of Slofstra who showed that the gapless problem $(1,1)$-gap is in general undecidable [Slo19]. Subsequently, Ji, Natarajan, Vidick, Wright and Yuen announced the major breakthrough result that MIP = RE, where RE is the class of recursively enumerable languages [JNV+20a]. More precisely, for any $10<s<$ 1 they construct a family of so-called synchronous games for which the $(1,s)$-gap problem decides the Halting problem. [^1] Note that for any $0<s<c\leq 1$, RE-hardness of the $(1,s)$-gap problem for a family of games immediately implies RE-hardness of the $(c,s)$-gap problem for the same family of games. We also remark that it was shown in [MNY20] that the complexity of the gapless problem is complete for the class $\Pi_{2}^{0}$, a class at the second level of the arithmetical hierarchy which contains both RE and coRE, i.e. there is a family of games for which the associated $(1,1)$-gap problem is $\Pi_{2}^{0}$-complete. This result indicates that, in general, approximating the quantum value of synchronous games within any additive constant is undecidable. However, this undecidability result does not imply that the quantum value is uncomputable for more specific classes of games. In fact, there are known classes of games for which the quantum value can be computed efficiently [Tsi87, Tsi93]. It is important to further explore which classes of nonlocal games allow for efficient computation or approximation of the quantum value. Indeed, understanding this could help identify scenarios where quantum entanglement offers an advantage over classical technologies and could lead to the development of new protocols for distributed tasks, such as delegating quantum computation [Gri19, KLVY23].
The general undecidability result of MIP=RE has also had significant mathematical consequences, most notably the resolution in the negative of the Connes embedding problem. Showing that it remains RE-hard to compute or approximate the quantum value of more specific classes of games could resolve other long-standing open problems as well, such as proving the existence of a non-hyperlinear group [PS23]. In fact, a similar undecidability result for so-called tailored non-local games recently led to the resolution of the Aldous–Lyons conjecture [BCLV24, BCV24].
In short, the complexity of approximating the quantum value of general nonlocal games has now been established. It is therefore the right time to embark on a research program to understand the following general yet more fine-grained questions, which motivate our work:
What is the complexity landscape of problems in $\mathrm{MIP}^{*}$? What is the complexity of the $(c,s)\textit{-}\mathrm{gap}^{*}$ problem for different natural classes of nonlocal games?
To answer the above fundamental questions, we must understand the following key points:
  1. I.
    What are natural well-structured classes of games which are MIP/RE-complete? [^2] This is analogous to NP-completeness in the classical setting. Recall that, say, 3SAT being NP-complete is a shorthand for: deciding satisfiability of 3SAT instances is NP-complete. Similarly a class of games is MIP-complete means: for some $0<s<c\leq 1$ deciding the $(c,s)$-gap problem for that class of games is MIP-complete. By MIP=RE [JNV+20a], this is equivalent to RE-completeness.
  2. II.
    How does the complexity of the $(c,s)$-gap problem depend on the class of games and the choice of the completeness and soundness parameters $c$ and $s$? In particular, is there a sharp transition from easy to undecidable in terms of the choice of $c,s$ for certain classes of games?
  3. III.
    How does limiting the question and/or answer size of a chosen family of games affect the complexity of the corresponding gapped promise problem?
There are two main approaches to addressing these questions: an algorithmic approach and a complexity-theoretic one. Our results focus on the latter. Before presenting them, we briefly review the current state of the art.
Algorithms. The first approach is to devise efficient algorithms to compute or approximate the quantum value of certain games, which is particularly relevant for question II. above. There are only a few examples of classes of games for which we have efficient algorithms to compute or approximate the quantum value. By a result due to Tsirelson [Tsi87, Tsi93], the quantum value of so-called XOR games can be computed efficiently by a semi-definite program (SDP). This work subsequently inspired the development of SDP-based approximation algorithms for several more general classes of games, including unique games and more recently 3-Coloring [KRT10, Bei10, CMS24]. So far, none of the existing approximation algorithms in the quantum setting have been shown to be optimal, in the sense that the existence of a better approximation algorithm would imply the collapse of certain complexity classes. In contrast, many classical approximation algorithms are known to be optimal unless P=NP. Showing quantum analogs of such results requires a better understanding of hardness reductions in the quantum setting, which we discuss next.
Complexity theory (hardness). The second approach is to study the hardness of gapped promise problems and to relate them to each other, which is relevant for all three questions I.III..
By [JNV+20a], synchronous games are MIP-complete in a strong sense: the $(c,s)$-gap for the class of synchronous games is MIP-complete for any $0<s<c\leq 1$. This result has since seen several refinements. Firstly, as mentioned above, it is now known that the same complexity theoretic result holds for a subclass of synchronous games called tailored games [BCV24]. Secondly, synchronous games with either constant question size [NZ23] or constant answer size [DFN+24] remain MIP-complete in a weaker sense [^3] These classes of games are MIP-complete in the following sense: for every choice of $0<s<c\leq 1$, there is a family of games with either constant question size or constant answer size such that the $(c,s)$-gap problem associated to it is MIP-complete. This does not mean that the same family of games (with fixed question size or fixed answer size) is MIP-complete in the strong sense. In fact, it follows from a simple argument that this cannot be the case, see the discussion in Remark 5.3., providing a partial answer to question III. above. Reducing the size of both the question and answer sets to at most a polynomial in the instance size is a major open question in the field, and is closely related to the quantum games PCP conjecture [NN24]. Both of the results [NZ23] and [DFN+24] are proven by reducing the Halting problem to a gapped promise problem for the class of games considered. To identify further MIP-complete classes of games, it is natural to consider reductions between gapped promise problems, instead of reductions from the Halting problem directly. This parallels the classical literature on hardness of approximation: for instance hardness of approximating 3SAT is proven by a reduction from another “stronger” hardness of approximation result, namely the PCP theorem [Hå01].
More precisely, a reduction from the $(c,s)$-gap problem for a family of games $C$ to the $(\overline{c},\overline{s})$-gap problem for a family of games $\overline{C}$ is a mapping $C\ni\mathcal{G}\mapsto\overline{\mathcal{G}}\in\overline{C}$ satisfying
  • (completeness) if $\omega^{*}(\mathcal{G})\geq c$, then $\omega^{*}(\overline{\mathcal{G}})\geq\overline{c}$, and
  • (soundness) if $\omega^{*}(\mathcal{G})<s$, then $\omega^{*}(\overline{\mathcal{G}})<\overline{s}$;
moreover this mapping should be efficient in an appropriate sense, see Section 2.3 for a formal definition. Crucially, such a reduction implies that if the $(c,s)$-gap problem for $C$ is hard, then the $(\overline{c},\overline{s})$-gap problem for $\overline{C}$ is also hard. To establish a reduction in the quantum setting, one usually needs to construct a quantum strategy for $\overline{\mathcal{G}}$ given a quantum strategy for $\mathcal{G}$ and vice-versa. This is greatly simplified when the given quantum strategy for $\mathcal{G}$, respectively $\overline{\mathcal{G}}$, is perfect. In this case, the operators from the given strategy will satisfy certain algebraic relations exactly (e.g. orthogonality or commutation relations) which can often be leveraged to directly construct a quantum strategy for $\overline{\mathcal{G}}$, respectively $\mathcal{G}$. Thanks to this simplification, several gapless reductions are known, meaning reductions between $(1,1)$-gap problems for various classes of games [Ji13, MRV15, Har24a, AKS19]. For instance, Ji showed that the $(1,1)$-gap problem for 3SAT games (in the clause-variable model) reduces to the $(1,1)$-gap problem for 3-Coloring games [Ji13], and Mančinska, Roberson, and Varvitsiotis showed that the $(1,1)$-gap problem for synchronous games reduces to the $(1,1)$-gap problem for independent set games, which we define shortly. On the other hand, when the starting strategy is not perfect, its operators will usually only satisfy certain algebraic relations approximately (e.g., they are only approximately orthogonal). In order to construct a quantum strategy for the target game, it is then often required to establish so-called stability theorems. Such theorems assert that if a set of operators approximately satisfies certain algebraic relations, then there exists a close-by set of operators that satisfies those relations exactly. For our application, the notion of closeness in a stability theorem must be well-behaved. More specifically, it should scale favorably with the number of operators involved – a property that is often highly non-trivial to establish. We refer to the discussion of our proof techniques below for further details. Such quantitative stability results usually lie at the heart of gap-preserving reductions in the quantum setting, i.e., reductions from a $(c,s)$-gap problem to a $(\overline{c},\overline{s})$-gap problem with $s<c$ and $\overline{s}<\overline{c}$.
Our main result shows how to lift the gapless reduction of Mančinska, Roberson, and Varvitsiotis to a gap-preserving reduction. To do so, we prove a new stability theorem, which we believe to be of independent interest, and which could be used to prove further gap-preserving reductions and elucidate the three questions I.III. discussed above.
Our results. The family of games we reduce to consists of independent set games. Given a graph $X$ and a positive integer $t$, the provers in the $t$-independent set game for $X$ aim to convince the verifier that the graph $X$ has an independent set of size $t$, i.e., a set of $t$ pairwise non-adjacent vertices. The question set $[t]=\{0,1,\ldots,t-1\}$ consists of $t$ labels, and the answer set $\mathcal{V}(X)$ consists of the vertices of $X$. To play the game, the verifier uniformly at random selects a pair $(i,j)$ from $[t]\times[t]$, and sends $i$ to the first prover and $j$ to the second. The provers then respond with vertices $u,v\in\mathcal{V}(X)$ and they win if and only if the following is satisfied:
  • if $i=j$, then $u=v$, and
  • if $i\neq j$, then $u$ and $v$ are distinct, non-adjacent vertices.
Our main technical result is the following.
Theorem A.
Let $\mathcal{G}$ be a synchronous game with uniform question distribution. Then there is a $t$-independent set game $\overline{\mathcal{G}}$, where $t$ is the number of questions of $\mathcal{G}$, such that the following hold:
  • if $\omega^{*}(\mathcal{G})=1$, then $\omega^{*}(\overline{\mathcal{G}})=1$, and
  • if $\omega^{*}(\mathcal{G})<1-\varepsilon$, then $\omega^{*}(\overline{\mathcal{G}})<1-\mathcal{O}(\frac{\varepsilon^{8}}{t^{25}})$.
As stated this theorem does not automatically give a gap-preserving reduction from any given class of synchronous games $C=\{\mathcal{G}_{x}\}_{x}$ to the class of independent set games, since the gap $\mathcal{O}(\frac{\varepsilon^{8}}{t_{x}^{25}})$ depends on the number of questions $t_{x}$ of the games $\mathcal{G}_{x}$ which could increase with $\left|x\right|$. However, as alluded to above, Natarajan and Zhang constructed in [NZ23] an RE-complete family of synchronous games with question sets of constant size, uniform question distribution, and with answer sets growing quasi-polynomially in $\left|x\right|$. Using this result, the following corollary then follows easily from Theorem A (see Section 5):
Corollary B.
There exist a constant $10<s<$ 1, a positive integer $t$, and a family $C$ of $t$-independent set games such that the $(1,s)$-gap problem for $C$ is $\mathrm{RE}$-complete.
Our results have several implications. First, they address the key points I.III. outlined above as follows:
  1. I.
    We show that independent set games are MIP/RE-complete. This class of games is a natural well-structured class of games which also plays a significant role in (classical) complexity theory.
  2. II.
    We demonstrate that there exists a constant $10<s<$ 1 for which the $(1,s)$-gap problem for independent set games is RE-hard and another constant where it becomes easy, in fact trivial, see Remark 5.3. However, the precise constant (if any) at which the problem transitions from RE-hard to easy remains unknown, see the future directions below for further details.
  3. III.
    Our MIP-complete class of games has constant question size, and is a more structured class than the class of synchronous games with constant question size from [NZ23]. This added structure opens the possibility of applying classically known techniques to compress the graphs of our independent set games, offering a new avenue for tackling the quantum games PCP conjecture.
Another intriguing implication arises from comparing our problem to its classical counterpart. The family of independent set games in Corollary B is succinctly presented, and the vertex sets (i.e., answer sets of the games) grow quasi-polynomially in $\left|x\right|$. By an exponential-time reduction, it follows that for the soundness parameter $10<s<$ 1 from Corollary B, the $(1,s)$-gap problem is undecidable for (non-succinctly presented) $t$-independent set games, whose vertex sets grow polynomially in $\left|x\right|$. [^4] By a similar exponential-time reduction NP equals NEXP. Since we are only asking whether the problem is decidable or not, we can consider such reductions, see also [CM24]. In contrast, the same problem is decidable in polynomial time in the classical setting. Indeed, given a graph with vertex set $\mathcal{V}$, it suffices to check if there is an independent set among the $t$-sized subsets of $\mathcal{V}$, and there are only $\binom{\left|\mathcal{V}\right|}{t}$ such subsets.
This is a striking phenomenon which adds to the list of quantum complexity-theoretic problems behaving entirely differently from their classical counterparts. Other examples include XOR games, which are NP-hard classically [Hå01], but as mentioned above, become easy in the quantum setting, and linear system games, for which the $(1,1)$-gap problem is efficiently decidable, but the $(1,1)$-gap problem is undecidable [Slo19]. Our results show an even more surprising phenomenon: let $t$ be as in Corollary B. Then for $t$-independent set games, the $(1,1)$-gap problem is efficiently decidable but the $(1,s)$-gap problem is undecidable for some constant $10<s<$ 1. This forms an additional motivation to study questions I.III., as it gives us a finer insight into the counterintuitive power of entanglement.
Finally, we remark that our proofs also apply to the commuting operator value instead of the quantum value, see Section 2.2 for a definition. There is a conjecture that MIPco=coRE. If this conjecture holds, our results would also imply that the same gapped promise problem as in Corollary B for the commuting operator value would be coRE-complete.
Proof techniques. Next, we briefly outline the main ingredients and the structure of the proof of our main Theorem A. Let $\mathcal{G}$ be a synchronous game as in the statement of the theorem. Let $Q$ denote its question set, $A$ its answer set, and $V$ its predicate, see Section 2.2 for definitions. The game graph $X(\mathcal{G})$ associated to $\mathcal{G}$ is defined as follows (see Definition 2.6 and [MRV15]): The vertices of $X(\mathcal{G})$ correspond to pairs $(q,a)\in Q\times A$, and two vertices $(q,a)$ and $(q^{\prime},a^{\prime})$ are adjacent if and only if $V(q,q^{\prime};a,a^{\prime})=0$ or $V(q^{\prime},q;a^{\prime},a)=0$. We then let $\overline{\mathcal{G}}$ in Theorem A be the $\left|Q\right|$-independent set game for the graph $X(\mathcal{G})$.
The proof now amounts to showing that the two bullet points in the statement of Theorem A are satisfied. The first bullet point is essentially taken care of by the gapless reduction of [MRV15], see Section 3.
The second bullet point is more involved, and we prove it in the contrapositive: Assuming we have a quantum strategy $\overline{\mathcal{S}}$ for the $\left|Q\right|$-independent set game for the graph $X(\mathcal{G})$ which wins with probability greater than $1-\mathcal{O}(\frac{\varepsilon^{8}}{t^{25}})$, we construct a strategy $\mathcal{S}$ which wins the synchronous game $\mathcal{G}$ with probability greater than $1-\varepsilon$. To construct the strategy $\mathcal{S}$, we need to take sums of operators from the strategy $\overline{\mathcal{S}}$. If the strategy $\overline{\mathcal{S}}$ were perfect, these sums of operators would directly form a valid quantum strategy. In particular, they would form a Projective Valued Measurement (PVM), i.e., a set of projections summing to the identity. However, as the strategy $\overline{\mathcal{S}}$ is only close to perfect, the resulting operators are only approximate projections which approximately sum to the identity. We thus need to round these sums of operators to actual PVMs, which we call rounded operators. [^5] In fact, for technical reasons, we actually round the sets of operators which we sum over and not the sums of operators. We refer to Section 5 for more details. This is where the following stability theorem comes into play, where for a tracial von Neumann algebra $(M,\tau)$, we define the two-norm $\left\lVert x\right\rVert_{2}^{2}\coloneqq\tau(x^{*}x)$ for $x\in M$; see Section 2.1 for more background on von Neumann algebras. If one is mainly interested in the quantum value, i.e., finite-dimensional strategies, then one can safely assume that $M$ is a matrix algebra with its usual (normalized) trace.
Theorem C.
[Corollary 4.2] Let $(M,\tau)$ be a tracial von Neumann algebra. Fix $10<\varepsilon<$ 1. Let $\{p_{j}\}_{1\leq j\leq m}$ in $M$ be projections which almost sum to the identity, i.e., $\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}\leq\varepsilon$. Then there exist projections $\{q_{j}\}_{1\leq j\leq m}$ in $M$ such that $\sum_{j=1}^{m}q_{j}=1$ and $\sum_{j=1}^{m}\left\lVert p_{j}-q_{j}\right\rVert_{2}^{2}\leq\mathcal{O}(m^{2}\varepsilon)$.
This stability theorem, and its variant for contractions Corollary A.1, improves on a previous result by Kim, Paulsen, and Schafhauser [KPS18, Lemma 3.5] (see also [Pad21, Har24b]), whose result gives an exponential dependence on $m$ for the closeness between the operators and the rounded operators.
The proof of Theorem C is partially inspired by de la Salle’s result [dlS22, Theorem 1.2], which is itself a strengthening and generalization of [JNV+20b, Theorem 5.2]). For the proof, we first show that we can perturb the projections $\{p_{j}\}_{j=1}^{m}$ so that $\sum_{j}E(p_{j})=1$, where $E:M\to\mathcal{Z}(M)$ denotes the conditional expectation of $M$ onto its center – for instance, if $M$ is a factor, then $E$ is nothing but the tracial state $\tau$. Subsequently, we use an amplification trick similar to [dlS22, Theorem 1.2], after which we have to do some additional effort to get the bounds from the statement. The reason for this additional effort, and the main difficulty in establishing Theorem C, is that we have to deal with the fact that the sum $\sum_{j}p_{j}$ may overshoot the identity (whereas [dlS22, Theorem 1.2] includes the assumption that this does not happen). Indeed, even if this sum is close to $1$ in $2$-norm, its operator norm could be as large as $m$, which is the reason for the dependence on $m$ in its conclusion. We also point out that the main difference between our approach and the approach from [KPS18] is that we perturb all projections at the same time, whereas the proof of [KPS18, Lemma 3.5] proceeds inductively, perturbing the projections one by one, yielding exponential dependence as compared to our polynomial factor of $m^{2}$.
Going back to the proof of our main Theorem A: in order to show that the PVM’s arising from the rounded operators from Theorem C give rise to a quantum strategy $\mathcal{S}$ for $\mathcal{G}$ with value greater than $1-\varepsilon$, we need a bit more. Indeed, the difference between the original operators and the rounded operators should not depend on both the number of questions and answers. However, the number of projections $m$ we are summing over in Theorem C, does depend on both the number of questions and answers.
Thus, we show a refinement of Theorem C where the difference between the original projections and the rounded ones only depends on the number of questions [^6] This is closely tied to the fact that for our application we rely on the result of [NZ23] as discussed above., see Theorem 4.1. In order to establish this finer result, we use the fact that some of the original projections in the strategy $\overline{\mathcal{S}}$ are already orthogonal. This gives an upper bound on the sum $\sum_{j}p_{j}$ which only depends on the number of questions, allowing us to finish the reduction and construct the desired strategy $\mathcal{S}$ for $\mathcal{G}$.
We note that the refined Theorem 4.1 is the cornerstone for our result and we are only able to show it thanks to our proof techniques described above. In particular, to the best of our knowledge, there does not seem to be a way to prove an analogous refinement by adapting the proofs of similar results from the literature, such as the aforementioned [KPS18, Lemma 3.5].
Related work. To the best of our knowledge, the only other known gap-preserving reductions in the quantum setting have been shown by Culf and Mastel in [CM24]. Our work and their work are complementary in many respects.
Stated in the language of this paper, Culf and Mastel show that for many NP-complete problems, there exists a constant $10<\overline{s}<$ 1 such that the associated $(1,\overline{s})$-gap problem is undecidable. For instance, they lift Ji’s gapless reduction for 3-Coloring [Ji13] to a gap-preserving reduction. In contrast, the problem we consider in this paper is a classically easy problem, which cannot be captured by their result.
In terms of techniques, Culf and Mastel’s work builds upon the framework developed by Mastel and Slofstra in [MS24]. [^7] In their work, Mastel and Slofstra perform a reduction from a gapped promise problem for synchronous games to a gapped promise problem for 3SAT albeit with a polynomially decreasing gap. Therefore, their techniques are -algebraic as they work at the level of the algebras associated to Boolean Constraint System games. Moreover, they consider problems with constant answer size and perform their gap-preserving reductions from the RE-complete family of synchronous games with constant answer size from [DFN+24]. In comparison, our techniques work at the level of the strategies, yielding a more direct approach to showing gap-preserving reductions. Furthermore, we consider a problem with constant question size, which cannot be captured by the techniques of Culf and Mastel, and we perform our gap-preserving reduction from the RE-complete family of synchronous games with constant question size from [NZ23]. Finally, we prove a new efficient stability theorem along the way, whereas Culf and Mastel rely on a known stability theorem from [CVY23] which doesn’t apply in our setting.
Another related work is that of Harris [Har24b] who analyzes Ji’s gapless reduction [Ji13] in the approximate regime and reaches a gap that decreases exponentially, which is not suitable for complexity-theoretic conclusions. We show in Appendix A that our stability result can be used to reduce the gap’s exponential decrease to a polynomial decrease. It is an interesting open question whether his work combined with our stability theorem could give a gap-preserving reduction.
Future directions. Our work opens many new avenues for future progress within our proposed research program guided by the three questions I.III. described above. Here we list a few specific open questions:
  1. 1.
    Complexity landscape of independent set games. In Corollary B, can we identify specific values of $t$ and $s$ for which the described problem is RE-hard? More precisely, how does the complexity of the $(1,s)$-gap problem for $t$-independent set games vary with $s$ for a fixed $t$? We show that it can be RE-hard as well as trivial, but could there be for instance a range of $s$ where it is complete for some other known complexity class?
  2. 2.
    Improving our stability theorem. Could the bound on the closeness between the operators and the rounded ones in Theorem C be improved? In particular, could it be possible to improve the bound so that it is independent of the number of operators $m$? [^8] For instance, could we improve the bound by considering different presentations of the corresponding -algebra? See also [CVY23] for similar questions for stability of groups. This would lead to a gap in Theorem A which is independent of the number of questions $t$. This could imply that, similarly to synchronous games, independent set games are MIP-complete in the strong sense. [^9] Recall that we say that a family of games is MIP-complete in the strong sense if the $(c,s)$-gap associated to it is MIP-complete for all $0<s<c\leq 1$. Note that here we refer to the general class of independent set games which do not have a fixed question size. Indeed, for any fixed $t\in\mathbb{N}$, the class of $t$-independent set games cannot be MIP-complete in the strong sense, see Remark 5.3.
  3. 3.
    Gapped reductions in a different regime. Beyond independent set games, it would be interesting to show reductions for $(c,s)$-gap problems where $c\neq 1$? So far all the known gap-preserving reductions rely on stability theorems which are tied to the regime where $c=1$. This question is also motivated from the theory of hardness of approximation in the classical setting where reductions are often performed for $(c,s)$-gap problems where $c\neq 1$.
Organisation of the paper. Besides the introduction, there are four other sections and an appendix. In Section 2 we collect some notation, definitions, and known results, and we describe the framework for our results. In Section 3 we give a streamlined proof of the gapless reduction from [MRV15], since we will use it for our main result. In Section 4 we prove our stability Theorem C and its refinement Theorem 4.1. Finally in Section 5 we complete the proof of our main results Theorem A and Corollary B. After the References, we include Appendix A where we collect some corollaries to the results from Section 4 for approximate POVMs and discuss their connections to the literature.
Acknowledgments. We thank Eric Culf, Kieran Mastel and Connor Paddock for helpful discussions. LM and TS are supported by the European Union via ERC grant QInteract (grant No 101078107) and VILLUM FONDEN via the QMATH Centre of Excellence (grant No 10059) and Villum Young Investigator (grant No 37532). PS is supported by an MSCA Fellowship (grant No 101111079) from the European Union.
2 Preliminaries
2.1 von Neumann algebras
A von Neumann algebra is a -subalgebra of the algebra $B(H)$ of bounded operators on some Hilbert space $H$ which is closed in the weak operator topology, i.e., the smallest topology such that all maps of the form $x\mapsto\langle x\xi,\eta\rangle$ for $\xi,\eta\in H$ are continuous. The famous von Neumann bicommutant theorem, see for instance [Tak01, Theorem II.3.9], states that $M\subset B(H)$ is a von Neumann algebra if and only if $M=M^{\prime\prime}$, where for a subset $S\subset B(H)$, $S^{\prime}\coloneq\{y\in B(H)\mid xy=yx\text{ for all }x\in S\}$ denotes the commutant of $S$. For simplicity, we will assume that our von Neumann algebras are represented on a separable Hilbert space $H$, i.e., $H$ has a finite or countable orthonormal basis.
We will be concerned with finite von Neumann algebras $M$, i.e., von Neumann algebras in which every isometry is a unitary. Equivalently, see for instance [Tak01, Theorem V.2.4], finite von Neumann algebras are exactly the von Neumann algebras which admit a positive linear functional $\tau:M\to\mathbb{C}$ with $\tau(1)=1$, which is
  • faithful, i.e., for all $x\in M$: $\tau(x^{*}x)=0\iff x=0$,
  • normal, i.e., its restriction to the unit ball of $M$ is continuous in the weak operator topology, and
  • tracial, i.e., for all $x,y\in M$: $\tau(xy)=\tau(yx)$.
We will refer to $\tau$ as a trace, and to the pair $(M,\tau)$ as a tracial von Neumann algebra. Given a tracial von Neumann algebra, we can associate a canonical trace-norm to it given by $\left\lVert x\right\rVert_{2}=\sqrt{\tau(x^{*}x)}$.
Example 2.1.
A first, and important, example of a tracial von Neumann algebra is the matrix algebra $\mathbb{M}_{n}(\mathbb{C})$ with its usual (normalized) trace. Readers interested solely in quantum strategies on finite-dimensional Hilbert spaces, can safely assume that all von Neumann algebras in the remainder of this paper are finite-dimensional, i.e., (direct sums of) matrix algebras.
2.2 Nonlocal games
In a (two-player) nonlocal game $\mathcal{G}$, two players, usually called Alice and Bob, play against a verifier. The game $\mathcal{G}$ is specified by four (finite) sets, consisting of question sets $Q$ and $Q^{\prime}$, and answer sets $A$ and $A^{\prime}$, together with a probability distribution $\pi$ on $Q\times Q^{\prime}$ and a predicate $V:Q\times Q^{\prime}\times A\times A^{\prime}\to\{0,1\}$. By enlarging the question and answer sets if necessary, we will assume for simplicity that the question set and answer set are the same for both players. In fact, we will mostly deal with synchronous games in this paper, where this assumption is enforced. We thus denote a nonlocal game $\mathcal{G}$ by a tuple $\mathcal{G}=(Q,A,\pi,V)$. During the game, the verifier samples a question pair $(q,q^{\prime})$ from $Q\times Q$ according to $\pi$, sends $q$ to Alice, $q^{\prime}$ to Bob, and they respond with answers $a,a^{\prime}\in A$ respectively. Alice and Bob win this round of the game if $V(q,q^{\prime};a,a^{\prime})=1$ and they lose otherwise. Importantly, while the setup of the game $\mathcal{G}=(Q,A,\pi,V)$ is known to everyone, Alice and Bob cannot communicate during the game, and thus have no knowledge of each other’s question and answer. Their goal is to come up with a strategy which maximizes their chances of winning.
Strategies for nonlocal games. A classical deterministic strategy for $\mathcal{G}$ consists of two functions $f:Q\to A$ and $f^{\prime}:Q\to A$ which Alice and Bob use to determine their answers. Using this strategy, the probability of winning the game $\mathcal{G}$ is
$\omega(\mathcal{G},\{f,f^{\prime}\})=\sum_{(q,q^{\prime})\in Q\times Q}\pi(q,q^{\prime})V(q,q^{\prime};f(q),f^{\prime}(q^{\prime})),$
and the classical value $\omega(\mathcal{G})$ (also denoted $\omega_{loc}(\mathcal{G})$) of the game $\mathcal{G}$ is the maximum of these probabilities over all possible such strategies. More generally, a classical strategy could also make use of shared randomness between the players. It is easy to see, however, that strategies using shared randomness do not provide an advantage over deterministic strategies.
On the other hand, a quantum strategy for $\mathcal{G}$ allows the players to determine their answers based on joint measurements on a shared finite-dimensional entangled state. This is modeled using so-called Positive Operator Valued Measurements (POVM), i.e., sets of positive operators on a Hilbert space which sum to the identity. However, it is well-known that by potentially enlarging the Hilbert spaces they act on, it is possible to consider Projective Value Measurements (PVMs) instead, i.e., sets of projections on a Hilbert space which sum to the identity. More precisely, a quantum strategy $\mathcal{S}$ for a nonlocal game $\mathcal{G}=(Q,A,\pi,V)$ consists of the following:
  1. 1.
    Two finite dimensional Hilbert spaces $H,H^{\prime}$;
  2. 2.
    For every $q\in Q$, a PVM $\{M_{q}^{a}\}_{a\in A}$ on $H$ and a PVM $\{N_{q}^{a}\}_{a\in A}$ on $H^{\prime}$;
  3. 3.
    A quantum state |$\ket{\psi}\in H\otimes H^{\prime}$.
Here we think of the $M_{q}^{a}$ as the operators from Alice’s measurement, and the $N_{q}^{a}$ as the operators from Bob’s measurement. Using this quantum strategy, the probability that the players measure and answer $a,a^{\prime}\in A$ when given the questions $q,q^{\prime}\in Q$ respectively is
$p(a,a^{\prime}\lvert q,q^{\prime})=\bra{\psi}M_{q}^{a}\otimes N_{q^{\prime}}^{a^{\prime}}\ket{\psi}.$
In particular, the probability of winning the game $\mathcal{G}$ using the strategy $\mathcal{S}$ is
$\omega^{*}(\mathcal{G},\mathcal{S})=\sum_{(q,q^{\prime})\in Q\times Q}\sum_{(a,a^{\prime})\in A\times A}\pi(q,q^{\prime})V(q,q^{\prime};a,a^{\prime})\bra{\psi}M_{q}^{a}\otimes N_{q^{\prime}}^{a^{\prime}}\ket{\psi},$
and the quantum value $\omega^{*}(\mathcal{G})$ of the game $\mathcal{G}$ is the supremum of $\omega^{*}(\mathcal{G},\mathcal{S})$ over all possible quantum strategies $\mathcal{S}$. We note that this quantum value is sometimes denoted by $\omega_{q}(\mathcal{G})$, but we stick to the above notation $\omega^{*}$ in this paper. We also point out that, unlike the case of classical strategies, there does not necessarily exist a quantum strategy realizing this supremum, see [Slo19].
One more common class of strategies for nonlocal games are the quantum commuting strategies. A quantum commuting strategy $\mathcal{S}$ for $\mathcal{G}=(Q,A,\pi,V)$ consists of the following:
  1. 1.
    A Hilbert space $H$;
  2. 2.
    For every $q\in Q$, two PVMs $\{M_{q}^{a}\}_{a\in A}$ and $\{N_{q}^{a}\}_{a\in A}$ on $H$ satisfying $M_{q}^{a}N_{q^{\prime}}^{a^{\prime}}=N_{q^{\prime}}^{a^{\prime}}M_{q}^{a}$ for all $q,q^{\prime}\in Q,a,a^{\prime}\in A$;
  3. 3.
    A quantum state |$\ket{\psi}\in H$.
One can again consider the probability that the players measure and answer $a,a^{\prime}\in A$ when given the questions $q,q^{\prime}\in Q$ respectively, which is given by $p(a,a^{\prime}\lvert q,q^{\prime})=\bra{\psi}M_{q}^{a}N_{q^{\prime}}^{a^{\prime}}\ket{\psi}$. The quantum commuting value $\omega_{qc}(\mathcal{G})$ of $\mathcal{G}$ is then the supremum of the winning probabilities over all quantum commuting strategies. Since we allow for arbitrary Hilbert spaces, it is not too difficult to show that this supremum is in fact attained by some quantum commuting strategy. Also, we note that the quantum commuting strategies contain the quantum strategies.
Synchronous games and synchronous strategies. A nonlocal game $\mathcal{G}$ is a synchronous game if it has just one question set $Q$ and one answer set $A$ for both Alice and Bob, and the predicate satisfies the following: for any $q\in Q$, $V(q,q;a,a^{\prime})=1$ if and only if $a=a^{\prime}$. In other words, given the same question, Alice and Bob are required to give the same answer in order to win the game.
Given a synchronous game $\mathcal{G}$, a strategy for $\mathcal{G}$ is called synchronous if the corresponding correlation satisfies $p(a,a^{\prime}\lvert q,q)=0$ if $a\neq a^{\prime}$ for any $q\in Q$. The synchronous quantum value $\omega^{*}_{s}(\mathcal{G})$ (or $\omega^{s}_{q}(\mathcal{G})$) of the synchronous game $\mathcal{G}$ is the supremum of the winning probabilities over all synchronous quantum strategies. Similarly, one can consider the synchronous quantum commuting value $\omega^{s}_{qc}(\mathcal{G})$ of the synchronous game $\mathcal{G}$, as the supremum of the winning probabilities over all synchronous quantum commuting strategies.
A natural question is whether one can characterize the synchronous strategies among general quantum (commuting) strategies. In fact, since different strategies can yield the same correlation, it is more natural to phrase this question in terms of the correlation sets $C_{q}$ and $C_{qc}$ which consist of the correlations $p(a,a^{\prime}\lvert q,q^{\prime})$ arising from quantum and quantum commuting strategies respectively, and their synchronous versions $C_{q}^{s}\subset C_{q}$ and $C_{qc}^{s}\subset C_{qc}$ for which we restrict to synchronous strategies. It turns out that there is an elegant answer established in [PSS+16], which we describe below. We note that the results in [PSS+16] are formulated in terms of $C^{*}$-algebras, whereas we choose to present them in terms of (tracial) von Neumann algebras, since this is more convenient for us in the remainder of the paper. Nevertheless, we point out that both formulations are equivalent in this case, merely by changing between the $C^{*}$-algebra and the von Neumann algebra generated by a given collection of PVMs.
Theorem 2.2 ([PSS+16, Corollary 5.6]).
A correlation $p(a,a^{\prime}\lvert q,q^{\prime})$ belongs to $C_{qc}^{s}$ (resp. $C_{q}^{s}$) if and only if there exists a tracial von Neumann algebra (resp. finite-dimensional von Neumann algebra) $M$ with trace $\tau$ and a generating family of projections $\{P_{q}^{a}\mid q\in Q,a\in A\}$ satisfying $\sum_{a\in A}P_{q}^{a}=1$ for every $q\in Q$ such that
$\forall q,q^{\prime}\in Q,a,a^{\prime}\in A:\quad p(a,a^{\prime}\lvert q,q^{\prime})=\tau(P_{q}^{a}P_{q^{\prime}}^{a^{\prime}}).$
Thanks to the above theorem, when computing the synchronous quantum (commuting) value of a game, one can work with collections of PVMs in tracial von Neumann algebras instead of dealing with actual strategies. We will mostly take this approach in the remainder of the paper, since a lot of our results boil down to perturbing the projections in approximate PVMs to get genuine PVMs.
Remark 2.3.
One can also characterize the correlations belonging to other correlation sets, such as $C_{loc}^{s}$ [PSS+16], and $C_{qa}^{s}$ and $C_{qs}^{s}$ [KPS18] in terms of traces. Our results can be interpreted accordingly for these correlations sets.
Given a synchronous game, another natural question to ask is how the quantum (commuting) value and the synchronous quantum (commuting) value are related. The following result due to de la Salle and Marrakchi [MdlS23] (and established by Vidick in the finite-dimensional setting in [Vid22]) is useful in translating between the two, see also [Pad25, Lin24]. In order to state the theorem formally, we call a probability distribution $\pi$ on $Q\times Q$ a $C$-diagonally dominant distribution if
$\pi(q,q)\geq C\cdot\max\left\{\sum_{q^{\prime}\in Q}\pi(q,q^{\prime}),\sum_{q^{\prime}\in Q}\pi(q^{\prime},q)\right\},$
for all $q\in Q$. For instance, if the distribution $\pi$ is uniform then we have $C=1/\left|Q\right|$ since:
$\pi(q,q)=\frac{1}{\left|Q\right|^{2}}=\frac{1}{\left|Q\right|}\sum_{q^{\prime}\in Q}\pi(q,q^{\prime})=\frac{1}{\left|Q\right|}\sum_{q^{\prime}\in Q}\pi(q^{\prime},q).$
Theorem 2.4 ([MdlS23]).
Suppose $\mathcal{G}$ is a synchronous game with a $C$-diagonally dominant question distribution. For $\mathrm{x}=q,qc$, if $\omega_{\mathrm{x}}(\mathcal{G})\geq 1-\varepsilon$, then $\omega^{s}_{\mathrm{x}}(\mathcal{G})\geq 1-\mathcal{O}(\left(\frac{\varepsilon}{C}\right)^{\frac{1}{4}})$.
Independent set games. Finally, we introduce the specific class of games we are mainly interested in, namely independent set games, which were first implicitly considered by Mančinska and Roberson in [MR14], and which were formally defined by Mančinska, Roberson, and Varvitsiotis in [MRV15]:
Definition 2.5.
Let $X$ be a simple graph with vertex set $\mathcal{V}(X)$ and let $t$ be a positive integer. Then the $t$-independent set game for $X$, denoted $(X,t)$, has question set $[t]=\{0,1,\ldots,t-1\}$, answer set $\mathcal{V}(X)$, and predicate $V:[t]\times[t]\times\mathcal{V}(X)\times\mathcal{V}(X)\to\{0,1\}$ satisfying $V(i,j;u,v)=1$ if and only if the following is satisfied:
  • if $i=j$, then $u=v$, and
  • if $i\neq j$, then $u$ and $v$ are distinct, non-adjacent vertices.
Every synchronous game has a game graph associated to it as in the following definition. Using this, we will later associate to a synchronous game $\mathcal{G}=(Q,A,\pi,V)$ the $\left|Q\right|$-independent set game for its game graph.
Definition 2.6 ([MRV15]).
Let $\mathcal{G}=(Q,A,\pi,V)$ be a synchronous game. The game graph of $\mathcal{G}$, denoted $X(\mathcal{G})$, is the undirected graph with vertex set $Q\times A$, and where $(q,a)$ is adjacent to $(q^{\prime},a^{\prime})$ if and only if $V(q,q^{\prime};a,a^{\prime})=0$ or $V(q^{\prime},q;a^{\prime},a)=0$.
For later reference, we explicitly compute the value of the $t$-independent set game for $X(\mathcal{G})$ when using a synchronous strategy. Let $\mathcal{S}$ be a synchronous strategy whose associated correlation is given by a collection of PVMs $\{P_{i}^{(q,a)}\}_{(q,a)\in\mathcal{V}(X)}$ indexed by $i\in[t]$ in a tracial von Neumann algebra $(M,\tau)$ as in Theorem 2.2. Then, it follows immediately from the definition that the losing probability of this strategy is
$\begin{split}1-\omega^{*}((X(\mathcal{G}),t),\mathcal{S})=\frac{1}{t^{2}}\sum_{\begin{subarray}{c}i\neq j,\\ (q,a)\end{subarray}}\tau\left(P_{i}^{(q,a)}P_{j}^{(q,a)}\right)+\frac{1}{t^{2}}\sum_{\begin{subarray}{c}i\neq j,\\ V(q,q^{\prime};a,a^{\prime})=0\,\mathrm{or}\\ V(q^{\prime},q;a^{\prime},a)=0\end{subarray}}\tau\left(P_{i}^{(q,a)}P_{j}^{(q^{\prime},a^{\prime})}\right).\end{split}$ (2.1)
2.3 Complexity theory
In this work, the computational problems we consider are so-called gapped promise problems.
Definition 2.7.
Let $0<s\leq 1$ be a constant. The gapped promise problem with completeness 1 and soundness $s$ for a family of games $C$ is the following problem. Given a nonlocal game $\mathcal{G}\in C$, decide whether:
  • $\omega^{*}(\mathcal{G})=1$ or
  • $\omega^{*}(\mathcal{G})<s$,
given the promise that one of the two cases holds. We denote this problem the $(1,s)$-gap problem for $C$.
The seminal result MIP=RE of Ji, Natarajan, Vidick, Wright, and Yuen [JNV+20b] shows that for any constant $0<s\leq 1$ there exists a family of games $C$ for which the $(1,s)$-gap problem is RE-complete, where RE is the class of recursively enumerable languages. The family of games they construct consists of synchronous games which are succinctly presented, see Definition 2.9.
Remark 2.8.
In this paper, we consider the question and answer sets of a game to be represented in binary. For a game $\mathcal{G}=(Q,A,\pi,V)$, we have $Q=\{0,1\}^{l}$ and $A=\{0,1\}^{m}$ for some $l,m\in\mathbb{N}$. We call $l,m$ the question size and the answer size of the game respectively.
Definition 2.9.
A family of games $C=\{\mathcal{G}_{n}\}_{n}$, where $\mathcal{G}_{n}=(Q_{n},A_{n},\pi_{n},V_{n})$, is said to be succinctly presented if there exist a probabilistic Turing machine $S$ and a Turing machine $V$ such that
  • on input $n$, $S$ samples a pair $(q,q^{\prime})\in Q\times Q$ with probability $\pi_{n}(q,q^{\prime})$,
  • on input $(n,q,q^{\prime},a,a^{\prime})\in\mathbb{N}\times Q\times Q\times A\times A$, $V$ outputs $V_{n}(q,q^{\prime};a,a^{\prime})$.
Furthermore, we call the succinctly presented family of games $C$ efficient if the question and answer sizes of $\mathcal{G}_{n}$ scale as $\mathcal{O}(\mathrm{poly}(n))$ and the above Turing machines $S,V$ can be chosen to run in time $\mathrm{poly}(n)$.
Remark 2.10.
In this paper, we only consider succinctly presented families of games which are also efficient. For the sake of brevity, we therefore omit the term “efficient” for succinctly presented families of games.
The reason for considering succinctly presented family of games is that we are interested in the complexity of computing the value of games where sampling the questions and verifying the answers can be done efficiently. We therefore consider the most general setting where the verifier in the game is efficient while the instance sizes may grow exponentially in the family parameter.
We can now formally state the result of MIP=RE as follows.
Theorem 2.11.
([JNV+20b]) For every constant $10<s<$ 1, there exists a succinctly presented family $C=\{\mathcal{G}_{n}\}_{n}$ of synchronous games whose question and answer sizes grow polynomially in $n$, such that the $(1,s)$-gap promise problem for $C$ is $\mathrm{RE}$-complete.
As discussed in the introduction, the above result is not sufficient for our application and we need the following refinement proved by Natarajan and Zhang in [NZ23].
Theorem 2.12.
([NZ23]) For every constant $10<s<$ 1, there exists a succinctly presented family $C=\{\mathcal{G}_{n}\}_{n}$ of synchronous games, with constant question size and with answer size scaling as $\mathrm{polylog}(n)$, such that the $(1,s)$-gap promise problem for $C$ is $\mathrm{RE}$-complete. Moreover, the question distribution of each game in $C$ is uniform.
Remark 2.13.
In Theorem 2.12, the smaller the soundness parameter $s$ is, the bigger the constant question size will be. Indeed, to decrease the soundness parameter, parallel repetition is used in [NZ23], which increases the question size. The exact dependence of the question size on the soundness parameter will not be necessary for our application.
We can now define gap-preserving reductions between promise problems. Intuitively, such a reduction consists of a map from instances of $C$ to instances of $\overline{C}$ which satisfies completeness and soundness properties. However, as it is important to work with succinctly presented problems, we need to phrase the definition formally using Turing machines which implement the map between instances.
Definition 2.14.
Let $0<s,\overline{s}\leq 1$ be constants, and let $C=\{\mathcal{G}_{n}\}_{n}$ and $\overline{C}=\{\overline{\mathcal{G}}_{m}\}_{m}$ be succinctly presented families of games where $\mathcal{G}_{n}=(Q_{n},A_{n},\pi_{n},V_{n})$ and $\overline{\mathcal{G}}_{m}=(\overline{Q}_{m},\overline{A}_{m},\overline{\pi}_{m},\overline{V}_{m})$. Let $S,V$ be any Turing machines associated to $C$ as in Definition 2.9. A reduction from the $(1,s)$-gap problem for $C$ to the $(1,\overline{s})$-gap problem for $\overline{C}$ consists of a probabilistic Turing machine $\overline{S}$ and Turing machine $\overline{V}$ satisfying the following conditions:
  • $\overline{S}$ has oracle access to $S$ and on input $n$ outputs $\overline{n}\in\mathbb{N}$ as well as $(\overline{q},\overline{q}^{\prime})\in\overline{Q}_{\overline{n}}\times\overline{Q}_{\overline{n}}$ with probability $\overline{\pi}_{\overline{n}}(\overline{q},\overline{q}^{\prime})$,
  • $\overline{V}$ has oracle access to $V$ and on input $(n,\overline{q},\overline{q}^{\prime},\overline{a},\overline{a}^{\prime})\in\mathbb{N}\times\overline{Q}_{\overline{n}}\times\overline{Q}_{\overline{n}}\times\overline{A}_{\overline{n}}\times\overline{A}_{\overline{n}}$ outputs $\overline{V}_{\overline{n}}(\overline{q},\overline{q}^{\prime},\overline{a},\overline{a}^{\prime})$,
  • $\overline{S}$ and $\overline{V}$ run in time $\mathrm{poly}(n)$,
  • (completeness) if $\omega^{*}(\mathcal{G}_{n})=1$, then $\omega^{*}(\overline{\mathcal{G}}_{\overline{n}})=1$,
  • (soundness) if $\omega^{*}(\mathcal{G}_{n})<s$, then $\omega^{*}(\overline{\mathcal{G}}_{\overline{n}})<\overline{s}$.
The intuition behind this definition of reduction is the following: a verifier playing a game $\mathcal{G}\in C$ could instead efficiently play the associated game $\overline{\mathcal{G}}\in\overline{C}$ as it can efficiently convert the Turing machines $S,V$ for $C$ to Turing machines $\overline{S},\overline{V}$ for $\overline{C}$. Due to the completeness and soundness properties, deciding the gapped promise problem for the family $\overline{C}$ must be as hard as deciding the gapped promise problem for the family $C$. In particular, we have the following:
Lemma 2.15.
If there is a reduction from the $(1,s)$-gap problem for $C$ to $(1,\overline{s})$-gap problem for $\overline{C}$ and deciding the $(1,s)$-gap problem for $C$ is $\mathrm{RE}$-hard, then the $(1,\overline{s})$-gap problem for $\overline{C}$ is also $\mathrm{RE}$-hard.
Proof.
This follows immediately from the completeness and soundness properties of a reduction. ∎
Remark 2.16.
A gapless reduction is a reduction as in Definition 2.14 where both $s=\overline{s}=1$. As mentioned in the introduction, there are a number of known examples of gapless reductions in the literature [Ji13, MRV15, Har24a, AKS19].
It is usually not too complicated to lift a gapless reduction to a reduction where the resulting gap $\overline{s}$ depends on the family parameter $n$, e.g. $s<1$ is constant but $\overline{s}=1-1/\mathcal{O}(\mathrm{poly}(n))$. For instance, this lifting can be done for the gapless reduction of [MRV15] in a relatively straightforward manner. However, such a reduction is not meaningful from a complexity theoretic point of view. Removing the dependence on $n$ from the gap is a challenging and subtle task, which is the main result of our work. Indeed, Theorem A can be seen as lifting the gapless reduction of [MRV15] to an actual gap-preserving reduction, where $10<s,\overline{s}<$ 1 are both constants.
3 Perfect regime
In this section, we review the proof of the following theorem due to Mančinska, Roberson, and Varvitsiotis [MRV15] in the framework from Section 2.2. Our main result Theorem A is a generalization of this result to the gapped case.
Theorem 3.1 ([MRV15]).
Let $\mathcal{G}=(Q,A,\pi,V)$ be a synchronous game, $X(\mathcal{G})$ be its corresponding game graph, and $t=\left|Q\right|$ be the number of questions. Then $\mathcal{G}$ has a perfect quantum strategy if and only if the independent set game $(X(\mathcal{G}),t)$ has a perfect quantum strategy..
Proof.
Since $\mathcal{G}$ is a synchronous game, any perfect strategy is necessarily synchronous. Hence, let $\mathcal{S}$ be a perfect synchronous strategy for the game $\mathcal{G}$ given by PVMs $\{P_{q}^{a}\}_{a\in A}$ for every $q\in Q$ on a tracial von Neumann algebra $(M,\tau)$. We can use this same strategy to play the independent set game $(X(\mathcal{G}),t)$. Indeed, as the independent set is of size $t=\left|Q\right|$, we can label questions in $[t]$ by elements $q\in Q$. When the provers receive question $q$ they measure the PVM $\{P_{q}^{a}\}_{a\in A}$ and answer with vertex $(q,a)\in Q\times A$ upon getting the measurement outcome $a\in A$. The losing probability of this strategy $\overline{\mathcal{S}}$ for the independent set game $(X(\mathcal{G}),t)$ is given by replacing the operators $P_{q}^{(q,a)}$ with $P_{q}^{a}$ in equation (2.1). With this replacement, the first sum in equation (2.1) is zero by definition and thus we get:
$0\displaystyle 1-\omega^{*}((X(\mathcal{G}),t),\overline{\mathcal{S}})=\frac{1}{t^{2}}\sum_{\begin{subarray}{c}q\neq q^{\prime},\\ V(q,q^{\prime};a,a^{\prime})=0\,\mathrm{or}\\ V(q^{\prime},q;a^{\prime},a)=0\end{subarray}}\tau(P_{q}^{a}P_{q^{\prime}}^{a^{\prime}})=$ 0 .
In other words, $\overline{\mathcal{S}}$ is a perfect strategy for $(X(\mathcal{G}),t)$. This concludes the first direction of the proof.
Conversely, let $\mathcal{S}^{\prime}$ be a perfect synchronous strategy for the independent set game $(X(\mathcal{G}),t)$ given by PVMs $\{P_{i}^{(q,a)}\}_{(q,a)\in Q\times A}$ for every $i\in[t]$ on a tracial von Neumann algebra $(M,\tau)$. Define the operators
$P_{q}^{a}\coloneqq\sum_{i}P_{i}^{(q,a)}.$
We show that for every $q\in Q$ the set of operators $\{P_{q}^{a}\}_{a\in A}$ forms a PVM. Firstly, we check that for all $q\in Q,a\in A$, the operators $P_{q}^{a}$ are projections. They are clearly self-adjoint, and furthermore
$P_{q}^{a}P_{q}^{a}=\sum_{i,j}P_{i}^{(q,a)}P_{j}^{(q,a)}=\sum_{i}P_{i}^{(q,a)}+\sum_{i\neq j}P_{i}^{(q,a)}P_{j}^{(q,a)}=\sum_{i}P_{i}^{(q,a)}=P_{q}^{a},$
where we use that since the strategy $\mathcal{S}^{\prime}$ is perfect, equation (2.1) implies that $P_{i}^{(q,a)}P_{j}^{(q,a)}=0$ for all $i\neq j$ [^10] We additionally make use of the fact that for $A,B$ positive, $AB=0$ if and only if $\tau(AB)=0$. The forward direction is clear. For the other direction, write $A=S^{*}S$, $B=T^{*}T$, then define $X=ST^{*}$ and notice that $0=\tau(AB)=\tau(X^{*}X)=\left\lVert X\right\rVert_{\tau}^{2}$ which implies that $X=0$ and thus $AB=S^{*}XT=0$. Next, we check that for every $q\in Q$, the operators $P_{q}^{a}$ and $P_{q}^{a^{\prime}}$ are orthogonal for $a\neq a^{\prime}$:
$P_{q}^{a}P_{q}^{a^{\prime}}=\sum_{i,j}P_{i}^{(q,a)}P_{j}^{(q,a^{\prime})}=\sum_{i}P_{i}^{(q,a)}P_{i}^{(q,a^{\prime})}+\sum_{i\neq j}P_{i}^{(q,a)}P_{j}^{(q,a^{\prime})}=0,$
where we use that $\{P_{i}^{(q,a)}\}_{(q,a)}$ forms a PVM to cancel the first sum. For the second sum, we notice that $V(q,q;a,a^{\prime})=0$ by synchronicity and thus we conclude from equation (2.1) as above that $P_{i}^{(q,a)}P_{j}^{(q,a^{\prime})}=0$ for all $i\neq j$. It remains to show that for every $q\in Q$ the operators in $\{P_{q}^{a}\}_{a\in A}$ sum to the identity. Note that as the operators in $\{P_{q}^{a}\}_{a\in A}$ are mutually orthogonal projections, we have $\sum_{a}P_{q}^{a}\leq 1$. Since $\{P_{i}^{(q,a)}\}_{(q,a)\in Q\times A}$ is a PVM for every $i$, we thus get
$\left|Q\right|=\sum_{i}\sum_{q,a}P_{i}^{(q,a)}=\sum_{q}\sum_{a}P_{q}^{a}\leq\sum_{q}1=\left|Q\right|,$
which implies that $\sum_{a}P_{q}^{a}=1$.
We now show that the synchronous strategy $\mathcal{S}$ given by the PVMs $\{P_{q}^{a}\}_{a\in A}$ for every $q\in Q$ on the tracial von Neumann algebra $(M,\tau)$ is a perfect strategy for the synchronous game $\mathcal{G}$. Indeed, the losing probability of the strategy $\mathcal{S}$ is given by
$1-\omega^{*}(\mathcal{G},\mathcal{S})$ $=\sum_{\begin{subarray}{c}q,q^{\prime},a,a^{\prime}\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\pi(q,q^{\prime})\tau(P_{q}^{a}P_{q^{\prime}}^{a^{\prime}})$
$=\sum_{\begin{subarray}{c}q,q^{\prime},a,a^{\prime}\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\sum_{i,j}\pi(q,q^{\prime})\tau(P_{i}^{(q,a)}P_{j}^{(q^{\prime},a^{\prime})})$
$=\sum_{i,j}\sum_{(q,a)\sim_{X(\mathcal{G})}(q^{\prime},a^{\prime})}\pi(q,q^{\prime})\tau(P_{i}^{(q,a)}P_{j}^{(q^{\prime},a^{\prime})})$
$\leq\sum_{i}\sum_{(q,a)\neq(q^{\prime},a^{\prime})}\pi(q,q^{\prime})\tau(P_{i}^{(q,a)}P_{i}^{(q^{\prime},a^{\prime})})$
$\qquad+\sum_{i\neq j}\sum_{(q,a)\sim_{X(\mathcal{G})}(q^{\prime},a^{\prime})}\pi(q,q^{\prime})\tau(P_{i}^{(q,a)}P_{j}^{(q^{\prime},a^{\prime})})$
$=0,$
where the last inequality follows from equation (2.1) together with the fact that the strategy $\mathcal{S}^{\prime}$ is perfect. Hence $\mathcal{S}$ is a perfect strategy for $\mathcal{G}$, finishing the proof of the theorem. ∎
Remark 3.2.
Since the quantum value of a game is not necessarily attained, the theorem as stated does not imply that $\omega^{*}(\mathcal{G})=1$ if and only if $\omega^{*}(X(\mathcal{G}),t)=1$. Nevertheless, one can prove that this is actually the case, by using that $\omega^{*}(\mathcal{G})=1$ if and only if for every $\varepsilon>0$, there exists a quantum strategy $\mathcal{S}$ for $\mathcal{G}$ with $\omega^{*}(\mathcal{G},\mathcal{S})>1-\varepsilon$. The “only if” of the above statement then follows in exactly the same way as in the proof of Theorem 3.1, since such a strategy $\mathcal{S}$ can be used for the independent set game $(X(\mathcal{G}),t)$ yielding quantum value $\omega^{*}((X(\mathcal{G}),t),\mathcal{S})>1-\varepsilon$. For the “if”-statement, one needs a stability theorem to round the projections of a given strategy for $(X(\mathcal{G}),t)$ in order to get a genuine strategy for $\mathcal{G}$. This is possible using stability theorems available in the literature, such as [KPS18, Lemma 3.5]. However, it also follows from our more general Theorem A, so we won’t go into the details here.
4 A stability theorem
In this section we prove our main stability result, Theorem C, and its more precise version Theorem 4.1. This result is the cornerstone for the proof of our main Theorem A. In order to make the results in this section applicable to both quantum strategies as well as quantum commuting strategies, the results and proofs are formulated in terms of a general tracial von Neumann algebra $(M,\tau)$, which we fix throughout this section. However, we point out that there is no harm in assuming $M$ to be finite-dimensional if one is mainly interested in quantum strategies.
The following is the main result of this section:
Theorem 4.1.
Fix $10<\varepsilon<$ 1. Assume $\{p_{i}^{a}\}_{1\leq i\leq t,1\leq a\leq s}$ is a family of projections in $(M,\tau)$ satisfying
  1. (1)
    for every $1\leq i\leq t$ and $1\leq a\neq a^{\prime}\leq s$: $p_{i}^{a}p_{i}^{a^{\prime}}=0$,
  2. (2)
    $\left\lVert\sum_{i,a}p_{i}^{a}-1\right\rVert_{2}\leq\varepsilon$.
Then there exist projections $\{q_{i}^{a}\}_{1\leq i\leq t,1\leq a\leq s}$ in $M$ such that $\sum_{i,a}q_{i}^{a}=1$ and $\sum_{i,a}\left\lVert p_{i}^{a}-q_{i}^{a}\right\rVert_{2}^{2}\leq 5t^{2}\varepsilon+\mathcal{O}(t^{3/2}\varepsilon^{3/2})$. In particular, we have
$\sum_{i,a}\left\lVert p_{i}^{a}-q_{i}^{a}\right\rVert_{2}^{2}\leq\mathcal{O}\left(t^{2}\left\lVert\sum_{i,a}p_{i}^{a}-1\right\rVert_{2}\right).$
Setting $s=1$, i.e., without the assumption that any of the original projections are orthogonal, we immediately get the following corollary, which is Theorem C from the introduction.
Corollary 4.2.
Let $\{p_{j}\}_{j=1}^{m}$ be projections in $(M,\tau)$ such that $\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}\leq\varepsilon<1$. Then there exist projections $\{q_{j}\}_{1\leq j\leq m}$ in $M$ such that $\sum_{j=1}^{m}q_{j}=1$ and $\sum_{j=1}^{m}\left\lVert p_{j}-q_{j}\right\rVert_{2}^{2}\leq 5m^{2}\varepsilon+\mathcal{O}(m^{3/2}\varepsilon^{3/2})$.
We also refer to Appendix A for some further corollaries to Theorem 4.1 concerning approximate POVMs. In particular, we answer a question from [Har24b] and generalize a result from [dlS22].
4.1 A technical lemma
Before starting the proof of Theorem 4.1, we prove Lemma 4.4 below, which, under the weaker assumption that only the center-valued traces of the rounded projections add up to 1, establishes the result with a sharper bound and rounded projections that commute with the original ones. It is in this lemma that the main von Neumann algebraic technicalities arise. Therefore, we first briefly recall the notions we need, and we refer to Example 4.3 below for a summary of the finite-dimensional case.
Given a tracial von Neumann algebra $(M,\tau)$, we denote its center by $\mathcal{Z}(M)\coloneqq\{y\in M\mid\forall x\in M:xy=yx\}$. We further denote by $E:M\to\mathcal{Z}(M)$ its center-valued trace, i.e., the unique trace-preserving normal conditional expectation from $M$ onto $\mathcal{Z}(M)$, see for instance [Tak01, Theorem V.2.6]. We say $M$ is a factor if $\mathcal{Z}(M)=\mathbb{C}1$. If $M$ is not a factor, then we can identify its center $\mathcal{Z}(M)$ with $L^{\infty}(X,\mu)$ for some probability space $(X,\mu)$, and we can canonically write $M$ as a direct integral of factors
[…]
see for instance [Tak01, Chapter IV.8]. In this setting, any element $a\in M$ can be written as a direct integral
[…]
Example 4.3.
If $M$ is a factor, then $E(x)$ is nothing but $\tau(x)1_{M}$. Furthermore, if $M$ is finite-dimensional, we do not need direct integrals. Indeed, any finite-dimensional tracial von Neumann algebra is a finite direct sum of matrix algebras $M=\oplus_{i=1}^{k}(\mathbb{M}_{n_{i}}(\mathbb{C}),\tau_{i})$, which is (4.1) above with $X=\{1,\ldots,k\}$, and where $\tau_{i}$ is the canonical trace on $\mathbb{M}_{n_{i}}(\mathbb{C})$ weighted in such a way that $\sum_{i=1}^{k}\tau_{i}(1)=1$. In this case, the decomposition of an element $a$ as in (4.2) is just the decomposition of the matrix $a$ into its block-diagonal form $a=\oplus_{i=1}^{k}a_{i}\in\oplus_{i=1}^{k}\mathbb{M}_{n_{i}}(\mathbb{C})$. In particular, if one is interested in the finite-dimensional case, the integrals at the end of the proof of Lemma 4.4 can be replaced by sums over the direct summands.
Lemma 4.4.
Let $\{p_{j}\}_{j=1}^{m}$ be projections in $(M,\tau)$. Then for $1\leq j\leq m$, there exist projections $r_{j}\in M$ commuting with $p_{j}$, such that $\sum_{j=1}^{m}E(r_{j})=1$ and $\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}\right\rVert_{2}^{2}\leq\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}$.
Proof.
First, assume $M$ is a type II1 factor. In particular, (1) $\mathcal{Z}(M)=\mathbb{C}1$, (2) for any $x\in M$, $E(x)=\tau(x)1$, and (3) for any projection $p\in M$ and any $t\in[0,\tau(p)]$, there exists a projection $q\leq p$ with $\tau(q)=t$. Set $\lambda\coloneq\left|\tau(\sum_{j=1}^{m}p_{j}-1)\right|$. We consider the following two cases:
  1. (1)
    $\tau(\sum_{j=1}^{m}p_{j})\leq 1$. In this case, we can choose for every $j\in\{1,\ldots,m\}$ a projection $\tilde{p}_{j}\leq 1-p_{j}$ such that $\tau(\tilde{p}_{j})=\frac{\lambda}{m}$. We then define $r_{j}\coloneq p_{j}+\tilde{p}_{j}$.
  2. (2)
    $\tau(\sum_{j=1}^{m}p_{j})>1$. In this case, we proceed inductively. [^11] Intuitively, we do the following: we have $m$ numbers between 0 and 1 adding up to more than 1. In order to make them sum to 1, we distribute the amount to be subtracted from each number as evenly as possible, while requiring each number to stay positive. Set $\lambda_{j}\coloneq\tau(p_{j})$ and define
    $\lambda^{(0)}_{j}$ :-$\coloneq\lambda_{j},\text{ and}:$
    $m^{(0)}$ :-#$\coloneq\#\{1\leq j\leq m\mid\lambda^{(0)}_{j}\neq 0\}.:$
    Now define inductively
    $\lambda^{(k+1)}_{j}$ :-$\coloneq\max\left\{0,\lambda^{(k)}_{j}-\frac{\sum_{l=1}^{m}\lambda_{l}^{(k)}-1}{m^{(k)}}\right\},\text{ and}:$
    $m^{(k+1)}$ :-#$\coloneq\#\{1\leq j\leq m\mid\lambda^{(k+1)}_{j}\neq 0\}.:$
    Since $m^{(0)}\geq m^{(1)}\geq\ldots\geq 0$, there is a first index $k\geq 0$ such that $m^{(k)}=m^{(k+1)}$. In this case, we see that for each $j\in\{1\leq j\leq m\mid\lambda^{(k)}_{j}\neq 0\}=\{1\leq j\leq m\mid\lambda^{(k+1)}_{j}\neq 0\}$:
    $\lambda^{(k+1)}_{j}=\lambda^{(k)}_{j}-\frac{\sum_{l=1}^{m}\lambda_{l}^{(k)}-1}{m^{(k)}}.$
    By construction, we thus have $0\leq\lambda^{(k+1)}_{j}\leq\lambda_{j}$, for $j\in\{1,\ldots,m\}$, satisfying
    $\sum_{j=1}^{m}\lambda^{(k+1)}_{j}=1.$
    Hence, we can choose a projection $\tilde{p}_{j}\leq p_{j}$ such that $\tau(\tilde{p}_{j})=\lambda_{j}-\lambda^{(k+1)}_{j}$. We now define $r_{j}\coloneq p_{j}-\tilde{p}_{j}$.
In both of the aforementioned cases, we see that
$\sum_{j=1}^{m}\tau(r_{j})=1,$ (4.3)
and
$\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}\right\rVert_{2}^{2}=\sum_{j=1}^{m}\tau(\tilde{p}_{j})=\left|\tau(\sum_{j=1}^{m}p_{j}-1)\right|\leq\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}.$ (4.4)
Next, assume $M$ is a type I factor, i.e., $M\cong\mathbb{M}_{n}(\mathbb{C})$ for some $n\in\mathbb{N}$. In particular, for every projection $p\in M$, we have $\tau(p)\in\{\frac{k}{n}\mid 0\leq k\leq n\}$, and hence $\lambda\coloneq\left|\tau(\sum_{j=1}^{m}p_{j}-1)\right|=\frac{k}{n}$ for some $k\in\mathbb{N}$. We can now perform the exact same construction as in the type II1 case, with the only additional requirement that we distribute $\lambda$ over the indices $j\in\{1,\ldots,m\}$ in multiples of $\frac{1}{n}$, thereby ensuring the existence of the projections $\tilde{p}_{j}$. We leave the details to the interested reader.
Finally, if $M$ is not a factor, we consider the canonical disintegration of $M$ as a direct integral of factors
[…]
over its center $\mathcal{Z}(M)$ identified with $L^{\infty}(X,\mu)$ for some probability space $(X,\mu)$. Writing, for $j\in\{1,\ldots,m\}$,
[…]
we can for every $x\in X$ repeat the above construction for $(p_{j,x})_{j=1}^{m}\in M_{x}$. Moreover, since $\lambda_{x}\coloneq\left|\tau(\sum_{j=1}^{m}p_{j,x}-1)\right|$ depends measurably on $x$, we can make the construction measurable, yielding projections
[…]
and by (4.4)
$\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}\right\rVert_{2}^{2}$ $=\sum_{j=1}^{m}\tau(\tilde{p}_{j}^{-}+\tilde{p}_{j}^{+})$
[…]
$=\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2},$
where we used Cauchy-Schwarz and the fact that $\mu$ is a probability measure in the last inequality. This finishes the proof of the lemma. ∎
4.2 Proof of Theorem 4.1
The remainder of the proof of Theorem 4.1 is inspired by [dlS22, Theorem 1.2]. Concretely, the first part of the following proof, notably the introduction of the matrix $x$, corresponds to the first half of the proof of [dlS22, Theorem 1.2]. However, in our case we do not have the guarantee that $\left\lVert x\right\rVert\leq 1$. Indeed, $\left\lVert x\right\rVert$ could potentially be as large as $t$ (see (4.10) in the proof), and therefore the estimates required for the upper bound appearing in the conclusion of Theorem 4.1 are quite different.
Proof of Theorem 4.1..
Let $r_{i}^{a}$ be the projections given by Lemma 4.4. Write $m=ts$, and fix an ordering of the pairs $j=(i,a)$, $1\leq j\leq m$. Consider the matrix
$x=\sum_{j=1}^{m}e_{j1}\otimes r_{j}p_{j}\in\mathbb{M}_{m}(\mathbb{C})\otimes M\cong\mathbb{M}_{m}(M),$
where $e_{ij}\in\mathbb{M}_{m}(\mathbb{C})$ denotes the elementary matrix with a 1 in the $ij^{\text{th}}$ entry and zeroes elsewhere. Note that $\mathbb{M}_{m}(\mathbb{C})\otimes M$ is a tracial von Neumann algebra, with trace given by $\operatorname{tr}_{m}\otimes\,\tau$, center $1_{m}\otimes\mathcal{Z}(M)$ and center-valued trace $E_{m}=\operatorname{tr}_{m}\otimes\,E$. In other words, given an $M$-valued matrix $(a_{ij})\in\mathbb{M}_{m}(M)$, we have $E_{m}((a_{ij}))=1_{m}\otimes(\frac{1}{m}\sum_{i=1}^{m}E(a_{ii}))$. Define the projections
$p\coloneq\sum_{j=1}^{m}e_{jj}\otimes r_{j}\in\mathbb{M}_{m}(M),\text{ and }q\coloneq e_{11}\otimes 1\in\mathbb{M}_{m}(M).$
Then $xq=px=x$, and $E_{m}(p)=1_{m}\otimes\frac{1}{m}\sum_{j=1}^{m}E(r_{j})=\frac{1}{m}=E_{m}(q)$. By a standard von Neumann algebra fact, see for instance [dlS22, Lemma 3.2], we can thus write $x=u\left|x\right|$ for a partial isometry $u$ satisfying $u^{*}u=q$ and $uu^{*}=p$. In particular, upon identifying $M\cong e_{11}\otimes M$ via $a\leftrightarrow e_{11}\otimes a$, the elements
$q_{j}\coloneq u^{*}(e_{jj}\otimes r_{j})u$
are projections in $M$ such that $\sum_{j=1}^{m}q_{j}=1$. Moreover, since $[p_{j},r_{j}]=0$, we have
$\left|x\right|q_{j}\left|x\right|=\left|x\right|u^{*}(e_{jj}\otimes r_{j})u\left|x\right|=x^{*}(e_{jj}\otimes r_{j})x=r_{j}p_{j}.$
Finally, we compute $\sum_{j=1}^{m}\left\lVert p_{j}-q_{j}\right\rVert_{2}^{2}$, and thereby show that the $q_{j}$, $j\in\{1,\ldots,m\}$, satisfy the conclusion of the theorem. First, note that we can write
$p_{j}-q_{j}=p_{j}-r_{j}p_{j}+(\left|x\right|-1)q_{j}\left|x\right|+q_{j}(\left|x\right|-1).$ (4.5)
We will bound the three terms individually. For the first term, we immediately get
$\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}p_{j}\right\rVert_{2}^{2}\leq\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}\right\rVert_{2}^{2}\leq\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}\leq\varepsilon.$ (4.6)
For the second and third term, more care is required, and since we will need to involve assumption (1) from the theorem, we will specifically write the indices $i$ and $a$ again.
First of all, we note that
$\left\lVert\sum_{j=1}^{m}r_{j}p_{j}-1\right\rVert_{2}$ $=\left\lVert\sum_{i,a}r_{i}^{a}p_{i}^{a}-1\right\rVert_{2}$
$\leq\left\lVert\sum_{i,a}(r_{i}^{a}p_{i}^{a}-p_{i}^{a})\right\rVert_{2}+\left\lVert\sum_{i,a}p_{i}^{a}-1\right\rVert_{2}$
$\leq\sum_{i}\left\lVert\sum_{a}(r_{i}^{a}p_{i}^{a}-p_{i}^{a})\right\rVert_{2}+\varepsilon$
$\leq\sqrt{t}\sqrt{\sum_{i}\left\lVert\sum_{a}(r_{i}^{a}p_{i}^{a}-p_{i}^{a})\right\rVert_{2}^{2}}+\varepsilon.$
Since for a fixed $i$, the projections $\{p_{i}^{a}\}_{a}$ are mutually orthogonal, we get by Pythagoras that
$\sqrt{t}\sqrt{\sum_{i}\left\lVert\sum_{a}(r_{i}^{a}p_{i}^{a}-p_{i}^{a})\right\rVert_{2}^{2}}$ $=\sqrt{t}\sqrt{\sum_{i,a}\left\lVert r_{i}^{a}p_{i}^{a}-p_{i}^{a}\right\rVert_{2}^{2}}$
$\leq\sqrt{t}\sqrt{\sum_{i,a}\left\lVert r_{i}^{a}-p_{i}^{a}\right\rVert_{2}^{2}}$
$\leq\sqrt{t}\sqrt{\varepsilon}.$
Altogether, we thus get
$\left\lVert\sum_{j=1}^{m}r_{j}p_{j}-1\right\rVert_{2}\leq\sqrt{t}\sqrt{\varepsilon}+\varepsilon.$ (4.7)
We will continue with bounding the third term in (4.5). Using the fact that $\sum_{j}q_{j}=1$, we get
$\sum_{j=1}^{m}\left\lVert q_{j}(\left|x\right|-1)\right\rVert_{2}^{2}$ $=\sum_{j=1}^{m}\tau((\left|x\right|-1)q_{j}(\left|x\right|-1))$
$=\tau((\sqrt{x^{*}x}-1)^{2})$
To get rid of the square root, we note that for positive real numbers $t$, we have the inequality $(\sqrt{t}-1)^{2}\leq(t-1)^{2}$. Through the continuous functional calculus for self-adjoint operators, we thus get the same inequality when replacing $t$ with any positive operator. Hence the above yields
$\begin{split}\sum_{j=1}^{m}\left\lVert q_{j}(\left|x\right|-1)\right\rVert_{2}^{2}&=\tau((\sqrt{x^{*}x}-1)^{2})\\ &\leq\tau((x^{*}x-1)^{2})\\ &=\tau\big{(}\big{(}\sum_{j=1}^{m}r_{j}p_{j}-1\big{)}^{2}\big{)}\\ &=\left\lVert\sum_{j=1}^{m}r_{j}p_{j}-1\right\rVert_{2}^{2}\\ &\leq t\varepsilon+2\sqrt{t}\varepsilon\sqrt{\varepsilon}+\varepsilon^{2},\end{split}$ (4.8)
where we used (4.7) for the last inequality. For the second term of (4.5), we compute similarly:
$\begin{split}\sum_{j=1}^{m}\left\lVert(\left|x\right|-1)q_{j}\left|x\right|\right\rVert_{2}^{2}&=\sum_{j=1}^{m}\tau(\left|x\right|q_{j}(\sqrt{x^{*}x}-1)^{2}q_{j}\left|x\right|)\\ &\leq\sum_{j=1}^{m}\tau(\left|x\right|q_{j}(x^{*}x-1)^{2}q_{j}\left|x\right|)\\ &=\sum_{j=1}^{m}\tau((x^{*}x-1)q_{j}\left|x\right|^{2}q_{j}(x^{*}x-1))\\ &\leq\left\lVert\sum_{j=1}^{m}q_{j}x^{*}xq_{j}\right\rVert\tau((x^{*}x-1)^{2})\end{split}$ (4.9)
Since the projections $q_{j}$ are mutually orthogonal, we have
$\left\lVert\sum_{j=1}^{m}q_{j}x^{*}xq_{j}\right\rVert\leq\max_{1\leq i\leq m}\left\lVert q_{i}x^{*}xq_{i}\right\rVert\leq\left\lVert x^{*}x\right\rVert.$
Moreover,
$\left\lVert x^{*}x\right\rVert=\left\lVert\sum_{j=1}^{m}r_{j}p_{j}\right\rVert=\left\lVert\sum_{i,a}r_{i}^{a}p_{i}^{a}\right\rVert\leq\sum_{i}\left\lVert\sum_{a}r_{i}^{a}p_{i}^{a}\right\rVert\leq t,$ (4.10)
where the last inequality holds since for every $i$, the element $\sum_{a}r_{i}^{a}p_{i}^{a}$ is a sum of mutually orthogonal projections, hence a projection. Combining this with (4.8) and (4.9), we thus get
$\sum_{j=1}^{m}\left\lVert(\left|x\right|-1)q_{j}\left|x\right|\right\rVert_{2}^{2}\leq t(t\varepsilon+2\sqrt{t}\varepsilon\sqrt{\varepsilon}+\varepsilon^{2}).$
Altogether, we conclude that
$\big{(}\sum_{j=1}^{m}\left\lVert p_{j}-q_{j}\right\rVert_{2}^{2}\big{)}^{\frac{1}{2}}$ $\leq\big{(}\sum_{j=1}^{m}\left\lVert p_{j}-r_{j}p_{j}\right\rVert_{2}^{2}\big{)}^{\frac{1}{2}}+\big{(}\sum_{j=1}^{m}\left\lVert(\left|x\right|-1)q_{j}\left|x\right|\right\rVert_{2}^{2}\big{)}^{\frac{1}{2}}+\big{(}\sum_{j=1}^{m}\left\lVert q_{j}(\left|x\right|-1)\right\rVert_{2}^{2}\big{)}^{\frac{1}{2}}$
$\leq\sqrt{\varepsilon}+\sqrt{t(t\varepsilon+2\sqrt{t}\varepsilon\sqrt{\varepsilon}+\varepsilon^{2})}+\sqrt{t\varepsilon+2\sqrt{t}\varepsilon\sqrt{\varepsilon}+\varepsilon^{2}}.$
Some straightforward calculations then give
$\sum_{i,a}\left\lVert p_{i}^{a}-q_{i}^{a}\right\rVert_{2}^{2}$ $=\sum_{j=1}^{m}\left\lVert p_{j}-q_{j}\right\rVert_{2}^{2}$
$\leq\varepsilon(t+\sqrt{t}+1+\sqrt{\varepsilon}(\sqrt{t}+1))^{2}$
$\leq 5t^{2}\varepsilon+O(t^{3/2}\varepsilon^{3/2}).$
This finishes the proof of Theorem 4.1. ∎
Remark 4.5.
Alternatively, using the inequality $(\sqrt{t}-1)^{2}\leq\left|t-1\right|=\sqrt{(t-1)^{2}}$ and Cauchy-Schwarz, we get in (4.8) the inequality
$\tau((\sqrt{x^{*}x}-1)^{2})\leq\tau(\sqrt{(x^{*}x-1)^{2}})\leq\left\lVert\sum_{j=1}^{m}r_{j}p_{j}-1\right\rVert_{2}\leq\sqrt{t}\sqrt{\varepsilon}+\varepsilon,$
and similarly in (4.9). Altogether, this leads to the bound
$\sum_{i,a}\left\lVert p_{i}^{a}-q_{i}^{a}\right\rVert_{2}^{2}\leq\mathcal{O}(t^{3/2}\varepsilon^{1/2}),$
which is better in $t$ but worse in $\varepsilon$. Since we will keep $t$ constant in the reduction below, we will use Theorem 4.1 as stated.
5 Approximate regime
In this section we will prove our main results, Theorem A and Corollary B. Firstly, using our stability Theorem 4.1, we will establish the soundness part of Theorem A, which together with the results from Section 3 will yield the desired result. Finally, we check that the construction yields a reduction in the sense of Definition 2.14, which then allows us to deduce Corollary B.
We start with the soundness part of Theorem A:
Lemma 5.1.
There exists a universal constant $\zeta>0$ such that the following holds. Let $\mathcal{G}=(Q,A,\pi,V)$ be a synchronous game with uniform question distribution, $X(\mathcal{G})$ be its corresponding game graph and $t=\left|Q\right|$ be the number of questions. For sufficiently small $\varepsilon>0$, if $\omega^{*}(\mathcal{G})<1-\varepsilon$, then $\omega^{*}((X(\mathcal{G}),t))<1-\zeta\frac{\varepsilon^{8}}{t^{25}}$.
Proof.
We prove the lemma in the contrapositive. Assume that the quantum value of the $t$-independent set game for $X(\mathcal{G})$ satisfies $\omega^{*}((X(\mathcal{G}),t))\geq 1-\zeta\frac{\varepsilon^{8}}{t^{25}}$ for some universal constant $\zeta$ to be determined later in the proof. Since the question distribution in the independent set game is uniform, Theorem 2.4 tells us that there is some universal constant $\xi$ such that the synchronous quantum value satisfies $\omega^{*}_{s}((X(\mathcal{G}),t))\geq 1-\xi\zeta\frac{\varepsilon^{2}}{t^{6}}$.
Following Theorem 2.2, let us consider a synchronous strategy $\overline{\mathcal{S}}$ corresponding to a family of PVMs $\{P_{i}^{(q,a)}\}_{(q,a)\in Q\times A}$ for $i\in[t]$ in a tracial von Neumann algebra $(M,\tau)$, which satisfies $\omega^{*}_{s}((X(\mathcal{G}),t),\overline{\mathcal{S}})\geq 1-\xi\zeta\frac{\varepsilon^{2}}{t^{6}}$. Note that we are technically not guaranteed that there exists such a strategy, only that there exist strategies $\overline{\mathcal{S}}_{\varepsilon^{\prime}}$ with $\omega^{*}_{s}((X(\mathcal{G}),t),\overline{\mathcal{S}}_{\varepsilon^{\prime}})>1-\xi\zeta\frac{\varepsilon^{2}}{t^{6}}-\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0$. However, doing the analysis below for $\overline{\mathcal{S}}_{\varepsilon^{\prime}}$ for a fixed $\varepsilon^{\prime}$ and then letting $\varepsilon^{\prime}\to 0$ would not change anything, so we omit this notational inconvenience.
Our goal is to construct a strategy $\mathcal{S}$ for $\mathcal{G}$ such that $\omega^{*}(\mathcal{G},\mathcal{S})\geq 1-\varepsilon$. In fact, the strategy $\mathcal{S}$ we construct is a synchronous strategy on the same von Neumann algebra $M$.
Let $\delta=1-\omega^{*}_{s}((X(\mathcal{G}),t),\overline{\mathcal{S}})$ be the losing probability of the strategy $\overline{\mathcal{S}}$. Consider for a fixed $q$ the set of operators $\{P_{i}^{(q,a)}\}_{(i,a)\in[t]\times A}$. This set of operators satisfies assumption (1) of Theorem 4.1 with $t=\left|Q\right|$ and $s=\left|A\right|$. Applying Theorem 4.1, we thus get for every $q\in Q$ a set of projections $\{Q_{i}^{(q,a)}\}_{(i,a)\in[t]\times A}$ such that $\sum_{i,a}Q_{i}^{(q,a)}=1$ and
$\sum_{i,a}\left\lVert P_{i}^{(q,a)}-Q_{i}^{(q,a)}\right\rVert_{2}^{2}\leq\kappa t^{2}\left\lVert\sum_{i,a}P_{i}^{(q,a)}-1\right\rVert_{2},$ (5.1)
for some universal constant $\kappa$. Moreover, for every fixed $q\in Q$, the operators $P_{i}^{(q,a)}$ satisfy
$\begin{split}\frac{1}{t^{2}}\sum_{(i,a)\neq(i^{\prime},a^{\prime})}\tau(P_{i}^{(q,a)}P_{i^{\prime}}^{(q,a^{\prime})})&=\frac{1}{t^{2}}\sum_{i,a\neq a^{\prime}}\tau(P_{i}^{(q,a)}P_{i}^{(q,a^{\prime})})+\frac{1}{t^{2}}\sum_{i\neq i^{\prime},a}\tau(P_{i}^{(q,a)}P_{i^{\prime}}^{(q,a)})\\ &\qquad\quad+\frac{1}{t^{2}}\sum_{i\neq i^{\prime},a\neq a^{\prime}}\tau(P_{i}^{(q,a)}P_{i^{\prime}}^{(q,a^{\prime})})\\ &\leq\delta,\end{split}$ (5.2)
where the first sum equals zero since $\overline{\mathcal{S}}$ is a synchronous strategy, and the bound on the other two terms follows from equation (2.1) by noting that $V(q,q;a,a^{\prime})=0$ for $a\neq a^{\prime}$ as the game is synchronous. In order to conclude that the projections $Q_{i}^{(q,a)}$ are close to the projections $P_{i}^{(q,a)}$, we now compute the following, where we denote by $\operatorname*{\mathbb{E}}_{q}\coloneqq\frac{1}{t}\sum_{q}$ the expectation over $q\in Q$:
$\operatorname*{\mathbb{E}}_{q}\left\lVert\sum_{i,a}P_{i}^{(q,a)}-1\right\rVert_{2}^{2}$ $=\operatorname*{\mathbb{E}}_{q}\tau\left(\sum_{i,i^{\prime},a,a^{\prime}}P_{i}^{(q,a)}P_{i^{\prime}}^{(q,a^{\prime})}+1-2\sum_{i,a}P_{i}^{(q,a)}\right)$
$=\operatorname*{\mathbb{E}}_{q}\tau\left(\sum_{(i,a)\neq(i^{\prime},a^{\prime})}P_{i}^{(q,a)}P_{i^{\prime}}^{(q,a^{\prime})}+1-\sum_{i,a}P_{i}^{(q,a)}\right)$
$\leq t^{2}\delta,$
where we used equation (5.2) together with the fact that
$\operatorname*{\mathbb{E}}_{q}\sum_{i,a}P_{i}^{(q,a)}=\frac{1}{t}\sum_{q,i,a}P_{i}^{(q,a)}=\frac{1}{t}\sum_{i}\sum_{q,a}P_{i}^{(q,a)}=1,$
since $\{P_{i}^{(q,a)}\}_{(q,a)\in Q\times A}$ is a PVM for every $i\in[t]$. Using Jensen’s inequality, we thus conclude that
$\operatorname*{\mathbb{E}}_{q}\left\lVert\sum_{i,a}P_{i}^{(q,a)}-1\right\rVert_{2}\leq\sqrt{\operatorname*{\mathbb{E}}_{q}\left\lVert\sum_{i,a}P_{i}^{(q,a)}-1\right\rVert_{2}^{2}}\leq t\sqrt{\delta}.$ (5.3)
Combining equation (5.1) and equation (5.3), we thus get
$\operatorname*{\mathbb{E}}_{q}\sum_{i,a}\left\lVert P_{i}^{(q,a)}-Q_{i}^{(q,a)}\right\rVert_{2}^{2}\leq\kappa t^{2}\operatorname*{\mathbb{E}}_{q}\left\lVert\sum_{i,a}P_{i}^{(q,a)}-1\right\rVert_{2}\leq\kappa t^{3}\sqrt{\delta}.$ (5.4)
In other words, the projections $Q_{i}^{(q,a)}$ are close “on average over $q\in Q$” to the projections $P_{i}^{(q,a)}$. We now construct the operators
$Q_{q}^{a}\coloneqq\sum_{i}Q_{i}^{(q,a)}.$
Since the projections $Q_{i}^{(q,a)}$ satisfy $\sum_{i,a}Q_{i}^{(q,a)}=1$ by construction, it follows immediately that the set $\{Q_{q}^{a}\}_{a\in A}$ is a PVM for every $q\in Q$.
By Theorem 2.2, we thus get a corresponding synchronous strategy $\mathcal{S}$ for the synchronous game $\mathcal{G}$ whose correlation satisfies
$p(a,a^{\prime}\lvert q,q^{\prime})=\tau(Q_{q}^{a}Q_{q^{\prime}}^{a^{\prime}}),$
for $q,q^{\prime}\in Q$ and $a,a^{\prime}\in A$. In particular, the losing probability for the synchronous game $\mathcal{G}$ incurred by using the synchronous strategy $\mathcal{S}$ is
$1-\omega^{*}(G,\mathcal{S})=\operatorname*{\mathbb{E}}_{q,q^{\prime}}\sum_{\begin{subarray}{c}a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\tau(Q_{q}^{a}Q_{q^{\prime}}^{a^{\prime}}).$ (5.5)
Note that by construction $\tau(Q_{q}^{a}Q_{q^{\prime}}^{a^{\prime}})=\sum_{i,i^{\prime}}\tau\left(Q_{i}^{(q,a)}Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)$. For the individual terms in this sum, we have
$\tau\left(Q_{i}^{(q,a)}Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)$ $=\left\lVert Q_{i}^{(q,a)}Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}$
$=\left\lVert Q_{i}^{(q,a)}\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)+\left(Q_{i}^{(q,a)}-P_{i}^{(q,a)}\right)P_{i^{\prime}}^{(q^{\prime},a^{\prime})}+P_{i}^{(q,a)}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}$
$\leq 4\left\lVert Q_{i}^{(q,a)}\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)\right\rVert_{2}^{2}$
$\qquad\quad+4\left\lVert\left(Q_{i}^{(q,a)}-P_{i}^{(q,a)}\right)P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}+4\left\lVert P_{i}^{(q,a)}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2},$
where we used that by the triangle inequality $\left\lVert x+y\right\rVert_{2}^{2}\leq 2\left\lVert x\right\rVert_{2}^{2}+2\left\lVert y\right\rVert_{2}^{2}$ for any operators $x,y$. Substituting this into equation (5.5), we get the following bound on the losing probability of the game $\mathcal{G}$ when using the strategy $\mathcal{S}$:
$1-\omega^{*}(G,\mathcal{S})$ $=\operatorname*{\mathbb{E}}_{q,q^{\prime}}\sum_{\begin{subarray}{c}a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\tau(Q_{q}^{a}Q_{q^{\prime}}^{a^{\prime}})$
$\leq\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,i^{\prime},q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\Big{(}\left\lVert Q_{i}^{(q,a)}\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)\right\rVert_{2}^{2}$
$\qquad\qquad\qquad\qquad\quad+\left\lVert\left(Q_{i}^{(q,a)}-P_{i}^{(q,a)}\right)P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}+\left\lVert P_{i}^{(q,a)}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}\Big{)}.$
We now consider each term separately. For the first term, we have
$\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,i^{\prime},q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\left\lVert Q_{i}^{(q,a)}\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)\right\rVert_{2}^{2}$ $\leq\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,i^{\prime},q,q^{\prime},a,a^{\prime}\end{subarray}}\tau\left(Q_{i}^{(q,a)}\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)^{2}\right)$
$=\frac{4}{t}\sum_{\begin{subarray}{c}i^{\prime},q^{\prime},a^{\prime}\end{subarray}}\tau\left(\left(Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)^{2}\right)$
$=\frac{4}{t}\sum_{\begin{subarray}{c}i^{\prime},q^{\prime},a^{\prime}\end{subarray}}\left\lVert Q_{i^{\prime}}^{(q^{\prime},a^{\prime})}-P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}$
$\leq 4\kappa t^{3}\sqrt{\delta},$
where the first inequality holds since we just sum over more positive elements, the second equality holds since $\sum_{i,q,a}Q_{i}^{(q,a)}=t$, and the last inequality follows by equation (5.4). Completely analogously, using the fact that $\sum_{i^{\prime},q^{\prime},a^{\prime}}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}=t$, we also get for the second term
$\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,i^{\prime},q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\left\lVert\left(Q_{i}^{(q,a)}-P_{i}^{(q,a)}\right)P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right\rVert_{2}^{2}\leq 4\kappa t^{3}\sqrt{\delta}.$
Finally, for the third term we get
$\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,i^{\prime},q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\tau\left(P_{i}^{(q,a)}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)$ $=\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i\neq i^{\prime},q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\tau\left(P_{i}^{(q,a)}P_{i^{\prime}}^{(q^{\prime},a^{\prime})}\right)$
$\qquad\qquad\qquad\qquad\quad+\frac{4}{t^{2}}\sum_{\begin{subarray}{c}i,q,q^{\prime},a,a^{\prime}:\\ V(q,q^{\prime};a,a^{\prime})=0\end{subarray}}\tau\left(P_{i}^{(q,a)}P_{i}^{(q^{\prime},a^{\prime})}\right)$
$\leq 4\delta+\frac{4}{t^{2}}\sum_{i,(q,a)\neq(q^{\prime},a^{\prime})}\tau\left(P_{i}^{(q,a)}P_{i}^{(q^{\prime},a^{\prime})}\right)$
$=4\delta$
where the inequality follows from applying equation (2.1) to the first sum, and observing that $V(q,q;a,a)=1$ by synchronicity, and thus $(q,a)\neq(q^{\prime},a^{\prime})$ in the second sum, and the last equality follows from the fact that $\{P_{i}^{(q,a)}\}_{(q,a)}$ is a PVM for every $i$.
Putting everything together, we conclude that $1-\omega^{*}(G,\mathcal{S})\leq 8\kappa t^{3}\sqrt{\delta}+4\delta\leq\kappa^{\prime}t^{3}\sqrt{\delta}$, for some universal constant $\kappa^{\prime}$ and thus by the definition of $\delta$,
$\omega^{*}(G,\mathcal{S})\geq 1-\kappa^{\prime}t^{3}\sqrt{\delta}\geq 1-\kappa^{\prime}\sqrt{\xi\zeta}\varepsilon.$
Hence defining the universal constant $\zeta$ as $\zeta=\frac{1}{\kappa\xi^{2}}$, we get that $\omega^{*}(G,\mathcal{S})\geq 1-\varepsilon$. This finishes the proof. ∎
We now have a complete proof of our main Theorem A, which we state here again for convenience:
See A
Proof.
Let $X(\mathcal{G})$ be the game graph of $\mathcal{G}$, and define $\overline{\mathcal{G}}$ to be the $t$-independent set game for $X(\mathcal{G})$. The first bullet point then follows from the “only if” direction of Theorem 3.1, see Remark 3.2. The second bullet point follows from Lemma 5.1. ∎
Next, we show that the construction from Theorem A yields a reduction in the sense of Definition 2.14.
Corollary 5.2.
Let $10<s<$ 1 be a fixed constant, let $C$ be the corresponding class of synchronous games from Theorem 2.12 and let $t$ be the associated constant number of questions in this family. Let $\overline{C}$ denote the class of $t$-independent set games. Then there exists a constant $10<\overline{s}<$ 1 such that the construction from Theorem A yields a reduction in the sense of Definition 2.14 from the $(1,s)$-$\mathrm{gap}^{*}$ problem for $C$ to the $(1,\overline{s})$-$\mathrm{gap}^{*}$ problem for $\overline{C}$.
Proof.
First, we construct an explicit map that takes instances of $C$ to instances of $\overline{C}$. Following Definition 2.14 of a reduction, we then show the existence of the required Turing machines which implement this map, as well as completeness and soundness of the reduction.
The map. We map a given synchronous game $\mathcal{G}_{n}\in C$ to $(X(\mathcal{G}_{n}),t)\in\overline{C}$ to the $t$-independent set game on the game graph $X(\mathcal{G}_{n})$ of $\mathcal{G}_{n}$.
Efficient Turing machines. Firstly, we note that the question size is constant for both the family $\{\mathcal{G}_{n}\}_{n}$ as well as the family $\{\overline{\mathcal{G}}_{m}\}_{m}$, and that their distributions are uniform. Therefore, we do not need to define Turing machines to sample questions.
Let $V$ be a Turing machine associated to $C$ for verifying question and answer pairs as in Definition 2.9. We now construct an efficient Turing machine $\overline{V}$ with oracle access to $V$, which checks the predicates of the $t$-independent set games for $X(\mathcal{G}_{n})$.
For $n\in\mathbb{N}$, questions $i,j\in[t]$ and answers $(q,a),(q^{\prime},a^{\prime})\in Q\times A$, the Turing machine $\overline{V}$ is defined as follows: it first compares the bits corresponding to $q$ and $q^{\prime}$ as well as $a$ and $a^{\prime}$. This is efficient as the number of bits is constant for the questions and grows as polylog$(n)$ for the answers. Then it uses its oracle access to $V$ in order to compute $V_{n}(q,q^{\prime};a,a^{\prime})$ and $V_{n}(q^{\prime},q;a^{\prime},a)$. The verifier accepts if and only if
  • $i=j$, $q=q^{\prime}$, $a=a^{\prime}$, or
  • $i\neq j$, $q\neq q^{\prime}$ and both $V_{n}(q,q^{\prime};a,a^{\prime})$ and $V_{n}(q^{\prime},q;a^{\prime},a)$ evaluate to one.
This is sufficient to cover all possible cases the verifier is supposed to check. Indeed, when $i=j$, the verifier accepts if and only if the answered vertices are the same. When $i\neq j$, the verifier accepts if and only if the answered vertices are distinct and non-adjacent. Checking that $q\neq q^{\prime}$ implies that the vertices are distinct. Moreover, they are non-adjacent if and only if both $V_{n}(a,a^{\prime}\lvert q,q^{\prime})$, $V_{n}(a^{\prime},a\lvert q^{\prime},q)$ evaluate to one. On the other hand, if $q=q^{\prime}$, the verifier immediately rejects. Indeed, in this case the vertices are either the same or adjacent, since whenever $V_{n}(a,a^{\prime}\lvert q,q)$, $V_{n}(a^{\prime},a\lvert q,q)$ evaluate to one it implies that $a=a^{\prime}$ by synchronicity.
Completeness. From Lemma 3.1 and Remark 3.2, it follows that for $\mathcal{G}_{n}\in C$, if $\omega^{*}(\mathcal{G}_{n})=1$, then $\omega^{*}((X(\mathcal{G}_{n}),t))=1$.
Soundness. Let the constant $\varepsilon$ be $\varepsilon=1-s$. From Lemma 5.1, if $\omega^{*}(\mathcal{G}_{n})<1-\varepsilon$ then $\omega^{*}((X(\mathcal{G}_{n}),t))<1-\mathcal{O}(\frac{\varepsilon^{8}}{t^{25}})$. Now let $\overline{s}$ equal $\overline{s}=1-\mathcal{O}(\frac{\varepsilon^{8}}{t^{25}})$. Then it is clear that $\overline{s}<1$. Moreover, $\overline{s}>0$ as we can always choose $t$ large enough, by adding dummy bits to the original questions.
This finishes the proof. ∎
We now have all the ingredients to deduce our main complexity-theoretic result Corollary B, which we state here again for convenience:
See B
Proof.
The result follows from combining Corollary 5.2, Theorem 2.12, and Lemma 2.15. ∎
We conclude with an important remark regarding the soundness parameter in Corollary B.
Remark 5.3.
For any fixed $t\in\mathbb{N}$, it is clear that there exists an $s$ for which the $(1,s)$-gap problem for $t$-independent set games becomes trivial as it is always possible to win a $t$-independent set game with some small probability. The provers could for instance decide on a set of $t$ vertices and answer according to that set of vertices. They would then at least win whenever the verifier asks them for the synchronous questions, i.e. when the questions are the same element in $[t]$. In fact, it is easy to see that this is a general phenomenon for any family of games with fixed question set or fixed answer set.
Appendix A Contractions which are almost projections
In this appendix, we describe some corollaries to Theorem 4.1 which involve perturbing approximate POVMs consisting of contractions which are almost projections to genuine PVMs. These results seem to be interesting in their own right, generalize some results from the literature, and answer an open question as described below. However, we collect them in this appendix, since they are not directly related to our main results.
The first corollary is relatively straightforward:
Corollary A.1.
Let $(M,\tau)$ be a tracial von Neumann algebra. Fix $10<\varepsilon,\delta<$ 1 and suppose $a_{1},\ldots,a_{m}\in M$ are positive contractions satisfying
  1. 1.
    $\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}\leq\varepsilon$, and
  2. 2.
    $\sum_{j=1}^{m}\left\lVert a_{j}^{2}-a_{j}\right\rVert_{2}\leq\delta$.
Then there exists a PVM $\{q_{j}\}_{1\leq j\leq m}$ in $M$ such that $\sum_{j=1}^{m}\left\lVert a_{j}-q_{j}\right\rVert_{2}^{2}\leq\mathcal{O}(m^{2}(\varepsilon+\delta))$.
Proof.
First of all, we use the well-known fact that for any positive contraction $a$, the spectral projection $p=1_{[1/2,1]}(a)$ satisfies
$\left\lVert p-a\right\rVert_{2}\leq 2\sqrt{2}\left\lVert a^{2}-a\right\rVert_{2},$
see for instance [KPS18, Lemma 3.4]. Applying this to the $a_{j}$’s, we get projections $\{p_{j}\}_{j=1}^{m}$ which by assumption (2) satisfy
$\sum_{j=1}^{m}\left\lVert p_{j}-a_{j}\right\rVert_{2}\leq 2\sqrt{2}\sum_{j=1}^{m}\left\lVert a_{j}^{2}-a_{j}\right\rVert_{2}\leq 2\sqrt{2}\delta.$
From assumption (1), we thus get
$\left\lVert\sum_{j=1}^{m}p_{j}-1\right\rVert_{2}\leq\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}+2\sqrt{2}\delta\leq\varepsilon+2\sqrt{2}\delta.$
The conclusion now follows from Corollary 4.2. ∎
Corollary A.1 yields polynomial dependence on $m$ in [Har24b, Theorem 5.13], thus answering [Har24b, Question 6.7]. It is not clear if one can use this (or a similar) result to eliminate the dependence on either $n$ or $m$ in that theorem.
It turns out that we can get the same conclusion as in Corollary A.1 under a much weaker assumption than (2) from Corollary A.1, see condition (2) below in Corollary A.2. This condition is the same as [dlS22, Theorem 1.2], and Corollary A.2 can be seen as a generalization of this theorem, where we don’t assume that $\{a_{j}\}_{j=1}^{m}$ is a POVM, but only an approximate POVM. Because of this, just like for Theorem 4.1, we get an additional factor of $m^{2}$ due to the fact that the operator norm of $\sum_{j}a_{j}$ my blow up.
Corollary A.2.
Let $(M,\tau)$ be a tracial von Neumann algebra. Fix $10<\varepsilon,\delta<$ 1 and suppose $a_{1},\ldots,a_{m}\in M$ are positive contractions satisfying
  1. 1.
    $\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}\leq\varepsilon$, and
  2. 2.
    $\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|\leq\delta$.
Then there exists a PVM $\{q_{j}\}_{1\leq j\leq m}$ in $M$ such that $\sum_{j=1}^{m}\left\lVert a_{j}-q_{j}\right\rVert_{2}^{2}\leq\mathcal{O}(m^{2}(\varepsilon+\delta))$.
Before we turn to the proof of Corollary A.2, we need an analogue of Lemma 4.4 for contractions, see Lemma A.4 below. We start with recording the following trivial observation which we will use repeatedly and without mention.
Lemma A.3.
Let $(M,\tau)$ be a tracial von Neumann algebra and $a\in M$ a positive contraction. Then for every $0\leq\lambda\leq\tau(a)$, there exists a positive element $0\leq b\leq a$ commuting with $a$ such that $\tau(b)=\lambda$.
Proof.
This follows immediately by letting $b=\frac{\lambda}{\tau(a)}a$. ∎
Lemma A.4.
Let $\{a_{j}\}_{j=1}^{m}$ be positive contractions in a tracial von Neumann algebra $(M,\tau)$. Then for $1\leq j\leq m$, there exist positive contractions $r_{j}\in M$ commuting with $a_{j}$ satisfying the following:
  1. 1.
    $\sum_{j=1}^{m}E(r_{j})=1$,
  2. 2.
    $\sum_{j=1}^{m}\left\lVert a_{j}-r_{j}\right\rVert_{2}^{2}\leq\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}$, and
  3. 3.
    $\left|\tau(\sum_{j=1}^{m}a_{j}r_{j}-1)\right|\leq\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|$.
Proof.
First, assume $M$ is a factor. We set $\lambda\coloneq\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|$ and distinguish two cases:
  1. 1.
    $\tau(\sum_{j=1}^{m}a_{j})\leq 1$. In this case, we can choose for every $j\in\{1,\ldots,m\}$ a positive contraction $\tilde{a}_{j}\leq 1-a_{j}$ commuting with $a_{j}$ such that $\tau(\tilde{a}_{j})=\frac{\lambda}{m}$. We then define $r_{j}\coloneq a_{j}+\tilde{a}_{j}$.
  2. 2.
    $\tau(\sum_{j=1}^{m}a_{j})>1$. In this case, exactly as in the proof of Lemma 4.4, we find numbers $0\leq\tilde{\lambda}_{j}\leq\tau(r_{j})$ such that $\sum_{j=1}^{m}\tilde{\lambda}_{j}=1$. We can thus find positive contractions $\tilde{a}_{j}\leq a_{j}$ commuting with $a_{j}$ such that $\tau(\tilde{a}_{j})=\tau(a_{j})-\tilde{\lambda}_{j}$. We define $r_{j}\coloneq a_{j}-\tilde{a}_{j}$.
In both cases, we see that the $r_{j}$’s are positive contractions satisfying
$\sum_{j=1}^{m}\tau(r_{j})=1,$
and, using that $x^{2}\leq x$ for $0\leq x\leq 1$,
$\sum_{j=1}^{m}\left\lVert a_{j}-r_{j}\right\rVert_{2}^{2}=\sum_{j=1}^{m}\left\lVert\tilde{a}_{j}\right\rVert_{2}^{2}=\sum_{j=1}^{m}\tau(\tilde{a}_{j}^{2})\leq\sum_{j=1}^{m}\tau(\tilde{a}_{j})=\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|\leq\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}.$
When $M$ is not a factor, we can again write $M$ as a direct integral over its center and proceed in exactly the same way as in the proof of Lemma 4.4. This yields (1) and (2) of the lemma. To check (3), we first note that
$\tau(\sum_{j=1}^{m}a_{j}\tilde{a}_{j})\leq\tau(\sum_{j=1}^{m}\tilde{a}_{j})=\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|.$
In the factorial case, we have
$\tau(\sum_{j=1}^{m}a_{j}r_{j}-1)$ $=\tau(\sum_{j=1}^{m}a_{j}(r_{j}-a_{j}))+\tau(\sum_{j=1}^{m}a_{j}^{2}-1)$
=±$=\pm\tau(\sum_{j=1}^{m}a_{j}\tilde{a}_{j})+\tau(\sum_{j=1}^{m}a_{j}^{2}-1),$
and thus
$\left|\tau(\sum_{j=1}^{m}a_{j}r_{j}-1)\right|\leq\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|.$
When $M$ is not a factor, then following the proof of Lemma 4.4, we see that
[…]
$=\tau((\sum_{j=1}^{m}a_{j}\tilde{a}_{j})1_{X^{-}})-\tau((\sum_{j=1}^{m}a_{j}\tilde{a}_{j})1_{X^{+}})+\tau(\sum_{j=1}^{m}a_{j}^{2}-1).$
In particular,
$\left|\tau(\sum_{j=1}^{m}a_{j}r_{j}-1)\right|$ $\leq\tau((\sum_{j=1}^{m}a_{j}\tilde{a}_{j})1_{X^{-}})+\tau((\sum_{j=1}^{m}a_{j}\tilde{a}_{j})1_{X^{+}})+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|$
$=\tau(\sum_{j=1}^{m}a_{j}\tilde{a}_{j})+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|$
$\leq\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|.$
This shows that (3) is satisfied and thus finishes the proof of the lemma. ∎
Proof of Corollary A.2.
Consider the subset $C\subset M^{m}$ given by
$C\coloneq\{(x_{1},\ldots,x_{m})\in M^{m}\mid\forall j:0\leq x_{j}\leq 1,x_{j}a_{j}=a_{j}x_{j},\sum_{j=1}^{m}E(x_{j})=1\}.$
Then $C$ is a weak-closed convex subset of the unit ball of $M^{m}$, hence compact. Moreover, by Lemma A.4, there exists a tuple $(r_{1},\ldots,r_{m})\in C$ satisfying $\left|\tau(\sum_{j=1}^{m}a_{j}r_{j}-1)\right|\leq\left|\tau(\sum_{j=1}^{m}a_{j}-1)\right|+\left|\tau(\sum_{j=1}^{m}a_{j}^{2}-1)\right|$. Consider the continuous affine map
$f:C\to\mathbb{C}:(x_{1},\ldots,x_{m})\mapsto\tau(\sum_{j=1}^{m}a_{j}x_{j}).$
Then $f$ attains its maximum at an extreme point, say $(p_{1},\ldots,p_{m})$. By the above, this maximum is larger than $f(r_{1},\ldots,r_{m})\geq 1-\varepsilon-\delta$. We also note that by construction,
$f(x_{1},\ldots,x_{m})=\tau(\sum_{j=1}^{m}a_{j}x_{j})\leq\tau(\sum_{j=1}^{m}x_{j})=\tau(\sum_{j=1}^{m}E(x_{j}))=1,$
for every $(x_{1},\ldots,x_{m})\in C$. Moreover, the proof of [dlS22, Lemma 3.1] implies that the extreme point $(p_{1},\ldots,p_{m})$ consists of projections.
The proof now proceeds similarly to the proofs of [dlS22, Theorem 1.2] and Theorem 4.1, but for the reader’s convenience, we sketch it below. Let
$x\coloneq\sum_{j=1}^{m}e_{j1}\otimes p_{j}a_{j}^{1/2}\in\mathbb{M}_{n}(M).$
Let $p=\sum_{j=1}^{m}e_{jj}\otimes p_{j}$ and $q=e_{11}\otimes 1$. Then, as before, we can write $x=u\left|x\right|$ with $uu^{*}=p$ and $u^{*}u=q$. We will again identify $M$ with $e_{11}\otimes M$ in the natural way. We now consider the projections
$q_{j}\coloneq u^{*}(e_{jj}\otimes r_{j})u\in M$
for $1\leq j\leq m$, and we note that $\sum_{j=1}^{m}q_{j}=1$, i.e., $\{q_{j}\}_{j=1}^{m}$ is a PVM. Furthermore,
$\left|x\right|q_{j}\left|x\right|=p_{j}a_{j},$
and as before we can decompose
$a_{j}-q_{j}=a_{j}-p_{j}a_{j}+(\left|x\right|-1)q_{j}\left|x\right|+q_{j}(\left|x\right|-1).$ (A.1)
For the first term, we immediately get
$\sum_{j=1}^{m}\left\lVert a_{j}-p_{j}a_{j}\right\rVert_{2}^{2}=\sum_{j=1}^{m}\tau((1-p_{j})a_{j}^{2})\leq\sum_{j=1}^{m}\tau((1-p_{j})a_{j})\leq 2\varepsilon+\delta.$
Before bounding the other two terms, we note that the same computation as for (4.7) together with the above gives
$\left\lVert\sum_{j=1}^{m}a_{j}p_{j}-1\right\rVert_{2}\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\left\lVert a_{j}p_{j}-a_{j}\right\rVert_{2}^{2}}+\left\lVert\sum_{j=1}^{m}a_{j}-1\right\rVert_{2}\leq\sqrt{m}\sqrt{2\varepsilon+\delta}+\varepsilon.$
Following (4.8) we thus get for the third term:
$\sum_{j=1}^{m}\left\lVert q_{j}(\left|x\right|-1)\right\rVert_{2}^{2}$ $=\tau((\sqrt{x^{*}x}-1)^{2})$
$\leq\tau((x^{*}x-1)^{2})$
$=\tau\big{(}\big{(}\sum_{j=1}^{m}a_{j}p_{j}-1\big{)}^{2}\big{)}$
$=\left\lVert\sum_{j=1}^{m}a_{j}p_{j}-1\right\rVert_{2}^{2}$
$\leq m(2\varepsilon+\delta)+2\sqrt{m}\varepsilon\sqrt{2\varepsilon+\delta}+\varepsilon^{2}.$
Finally, following (4.9), (4.10), and the above, we get for the second term:
$\sum_{j=1}^{m}\left\lVert(\left|x\right|-1)q_{j}\left|x\right|\right\rVert_{2}^{2}$ $\leq\left\lVert\sum_{j=1}^{m}q_{j}x^{*}xq_{j}\right\rVert\tau((x^{*}x-1)^{2})$
$\leq m(m(2\varepsilon+\delta)+2\sqrt{m}\varepsilon\sqrt{2\varepsilon+\delta}+\varepsilon^{2}).$
The proof now finishes in exactly the same way as the proof of Theorem 4.1. ∎